### v2

# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

Abstract
The accomplished functions of a variety of quantum devices are closely associated with the controlling of exciton transport. To this end we study the exciton transport of the two-dimensional system consisting of two-level multichains with various coupling configurations in a cavity. Two types of the chains are considered, including the Tavis-Cummings and the Su-Schrieffer-Heeger chain. Two inter-chain coupling conformations are considered, including square and triangle type. The effects of the inter-chain coupling, the dimerization parameter, the cavity, the length and number of the chains on the exciton transport are in detail investigated for different coupling configurations through the spectra and steady-state dynamics of the system. The results show that in the absence of cavity the exciton transport currents and efficiency is decided by the distribution of population of exciton on the multichains. However, when the cavity is considered the exciton transport currents and efficiency is controlled by the cavity-dressed energy level crossings and anticrossings near the zero-energy modes, at which the coherent excitation and Landau-Zener transitions occur. Therefore, the exciton transport can be enhanced or suppressed at the crossings and anticrossings, in which the polariton acts as crucial role. Besides, it is discovered that the exciton transport efficiency is closely related with the parity of both the length and number of the chains. This work is important for the understanding of the exciton transport mechanism in the multichain-cavity system, and provides theoretical basis for excitonic devices with controllable and efficient exciton transport.
I Introduction
In quantum device, the carriers of energy and information transport are mainly electron, photon, and exciton. Enhancing and controlling the efficiency, the speed, the distance and the current of carrier transport are the center to accomplish the functions of various quantum devices, such as organic solar cells wurfel2015impact ; lu2022simultaneously ; melianas2017charge , organic light-emitting diodes shan2011transport ; tanase2003unification ; shen2022through , quantum dot solids zhang2022ultrafast , organic semiconductor devices sneyd2021efficient , quantum simulation, and quantum computation georgescu2014quantum ; yang2019quantum etc. The exciton loss caused by different factors, including the dissipation cainelli2021exciton ; zhu2024quantum ; bose2022effect , exciton-exciton annihilation wittmann2020energy , disorder zhu2022single ; quenzel2022plasmon ; schachenmayer2015cavity and defect sites wagner2011bound ; akselrod2014visualization , is one of the influence on the efficiency in a long transport distance. Thus, a large of studies are devoted to revealing the exciton transport mechanisms and properties of various quantum system engel2007evidence ; menke2013tailored , and to achieving the requirements for the property of exciton transport through reduce and avoid detrimental factors.
In experiment, the one-dimensional (1D) system of exciton or photon transport can be realized by various platforms, including organic aggregates molecule chainssu1979solitons , superconducting qubit circuits roushan2017spectroscopic ; xu2018emulating ; blais2003tunable , Rydberg-atom chains barredo2015coherent ; orioli2018relaxation , plasmons su2021perovskite ; seitz2020exciton , semiconductorssneyd2021efficient ; giannini2022exciton ; muller2023directed . At room temperature the strong exciton-photon coupling between organic aggregates and microcavities lidzey1998strong ; lidzey1999room ; kena2008strong ; tischler2005strong ; coles2014polariton or plasmons bellessa2014strong ; bellessa2004strong ; zengin2015realizing ; li2018plasmon can result in new applications like polariton lasing kena2010room , in which the molecular energy structures are modified galego2015cavity and the polaritons form between the edge state excitons and the cavity photons kavokin2003cavity ; agranovich2003cavity . The presence of polaritons can enhance the efficiency wei2019enhanced ; wei2022cavity , rate wang2021polariton and distance du2018theory ; myers2018polariton of the exciton transport because the decay of excitons on chains is suppressed.
The hybrid system consisting of an 1D molecule chain and a cavity can be described by the Tavis-Cummings (TC) model tavis1968exact for uniform coupling and the Su-Schrieffer-Heeger (SSH) model su1979solitons for the staggered hopping interaction. Based on the TC model, the interaction strength of the collective $N-$atom can be predicted fink2009dressed , a new type of quantum phase transitions was discovered at quite low coupling strengthszou2013quantum . Specially, the study shows that the SSH chain can be tuned into topologically trivial and nontrivial phases for the even number of the two-level systems (TLSs) cai2019observation . For the even chain, in the topologically nontrivial regime, a small gap bewteen two localized nearly-zero-energy modes arises in the middle of the topological bandgap, which leads to the formation of the hybrid edge modes with the left and the right edge state. Therefore, the nontrivial edge states topologically protected can enhance the exciton transport for a finite-size SSH chain, which is verified by the theoretical and experimental results chang2020probing ; wei2022cavity . In fact, the nature of the gap is the energy anticrossing, at which the Landau–Zener (LZ) transitions occur. In the two-dimensional (2D) and three-dimensional tungsten disulfide system, the anticrossing dispersions were presented, at which the boosting transport capability of excitons was corroborated in theoretically and experimentally guo2022boosting ; yuan2017exciton . Thus, the properties of the more than 1D system are fascinating and valuable to study on the exciton transport.
Motivated by the works on the high dimensional tungsten disulfide, we devote to the studies of the joint effect of inter-chain on exciton transport in a multi-chain TLSs with cavity in this paper. Each chain contains two alternative types, i.e., the TC and the SSH chain, and two conformations of the inter-chain coupling are set, i.e., square and triangle type. Firstly, in order to reveal the mechanism of exciton transport dynamics of the hybrid system in the condition of the strong exciton-cavity coupling, the energy spectrum with the inter-chain coupling strength and the dimerization parameter is presented through obtaining the eigenvalues of the Hamiltonian for six configurations. Meanwhile, we calculate the von Neumann entropy of exciton and Hopfield coefficients of photon to confirm the locations of the crossings and anticrossings of the near-zero energy mode dressed by the cavity. What’s more, the effects of the inter-chain coupling, the dimerization parameter, the cavity, the TLS transition frequency detuning, the number and length of the chains on the steady-state dynamics of exciton transport is investigated for different configurations by numerically solving the Lindblad quantum master equation under the TLS drive and TLS+cavity drive. In the absence of cavity, the channel of exciton transport only is the multichains, in which the efficiency depends on the distribution of exciton on the whole multichains. Importantly, for the hybrid system the exciton transport is obviously related to the crossings and anticrossings of the dressed near-zero energy modes due to the coherent excitation and LZ transition between the multichains and the cavity, at which the excitation is exchanged. As a consequence the exciton transport is controlled by the crossings and anticrossings, in which the enhancing or suppressing effects is decided by the configuration of the multichains and the polaritons act as important role. The optimal parameters regions on the space of the TLS transition frequency detuning, the inter-chain coupling strength, and the dimerization parameter are presented. Further, the results show that the exciton transport dynamics depend on the odevity of the number and length of the chains, i.e. the efficiency behaves oscillation with them.
The remainder of this paper is organised as follows: In Sec. II the theoretical model is introduced including the Hamiltonian of the multichain-cavity coupling system (in Sec. II.1), the formula calculating the von Neumann entropy of excitons (in Sec. II.2), and the Lindblad quantum master equation of the system with external driving field (in Sec. II.3). In Sec. II.4 the exciton transport currents and efficiency for the steady state are derived. In Sec. III the energy spectrums are given by numerically diagonalizing the Hamiltonian. In Sec. IV we present and discuss the steady-state dynamics of excitons for various configurations of the multichains, in which the effects of the cavity, the inter-chain coupling, the dimerization parameter, the TLS transition frequency detuning (in Sec. IV.1), the number and length of the chains (in Sec. IV.2) on exciton transport are analysed. Section VI is a brief summary and future perspective.
Figure 1: (color online) The schematic representation of the hybrid system that the 2D multi-chain TLSs are placed in a cavity for the TC ST (a), TC TT (b), and SSH ST (c). The red sphere represents the TLS with ground state $|g\rangle$ and excited state $|e\rangle$ between which the transition frequency is $\omega_{\rm a}$ and the decay rate is $\gamma$. The decay rate of the imperfect cavity is $\kappa$. The 2D TLSs array can be driven by an external coherent field with driving frequency $\omega$ via the first TLS of the first row with driving amplitude $\xi_{{\rm T},1}$ or/and the cavity with driving amplitude $\xi_{c}$. The excitons in the multichains finally outflow from the last column into a reservoir (drain) with rate $\gamma_{\rm R}$. For the TC chains and SSH chains, the intra-chain coupling strengths are respectively denoted by the $h_{\rm TC}$ and $h_{\rm SSH}$. $\delta$ is the dimerization strength of SSH chain. It exists the inter-chains couplings with the strength $\zeta$ between the nearest TLSs in two nearest chains.
II Model
II.1 Hamiltonian
The considered system is a 2D molecular aggregate placed in a single-mode cavity, which can be modelled as multichains with $L$ ($L\geqslant 2$) chains as shown in Fig. 1. The $k$th chain has $N_{k}$ identical TLSs. For simplicity, we set that the numbers of the TLSs for different chains are the same ($N_{1}=N_{2}=...=N_{L}=N$). The Hamiltonian of the multichain-cavity coupling system can be expressed as
$H_{0}=H_{\rm e}+H_{\rm c}+H_{\rm ec}$ (1)
The $H_{\rm e}$ represents the Hamiltonian of the multichains, and contains two terms as
$H_{\rm e}=\sum^{L}_{k=1}H_{{\rm e},k}+H_{\rm e,inter-chain}$ (2)
Here, $H_{{\rm e},k}$ are the Hamiltonian of the $k$th bare chain. Each chain can be set as different models.
For the uniform coupling strength between two nearest-neighbor TLS monomers in a chain, under the rotating frame of the driving frequency $\omega$ the Hamiltonian of the $k$th chain is described by the TC model as
$H_{{\rm e},k}=$ $-\sum_{i=1}^{N}\Delta_{\rm a}\sigma_{k,i}^{+}\sigma_{k,i}^{-}-$
$h_{\rm TC}\sum_{i=1}^{N-1}(\sigma_{k,i}^{+}\sigma_{k,i+1}^{-}+{\rm H.c.}),$ (3)
where $h_{\rm TC}$ characterizes the coupling strength between two nearest-neighbor TLS monomers in a chain, $\sigma_{k,i}^{+}=|e_{k,i}\rangle\langle g_{k,i}|$ ($\sigma_{k,i}^{-}=|g_{k,i}\rangle\langle e_{k,i}|$) is the raising (lowering) operator of the TLS with same energy $\omega_{\rm a}$ because of the identical partical (setting $\hbar=1$) at the $i$th site of the $k$th chain with the groud state $|g_{k,i}\rangle$ and the excited state $|e_{k,i}\rangle$, $\Delta_{\rm a}=\omega-\omega_{\rm a}$ is the TLSs transition frequency detuning relative to driving field frequency $\omega$, ${\rm H.c.}$ denotes the Hermitian conjugate.
Moreover, for the staggered hopping interaction between the nearest-neighbor TLS monomers in a chain, the Hamiltonian of each chain is described by the SSH model as
$H_{{\rm e},k}=$ $-\sum_{i=1}^{N}\Delta_{\rm a}\sigma_{k,i}^{+}\sigma_{k,i}^{-}-h_{\rm SSH}\times$
$\sum_{i=1}^{N-1}\left[1+(-1)^{i}\delta\right](\sigma_{k,i}^{+}\sigma_{k,i+1}^{-}+{\rm H.c.}),$ (4)
where $h_{\rm SSH}$ is similar to the $h_{\rm TC}$, and $\delta$ represents the dimerization strength.
The inter-chain coupling term is denoted by the $H_{\rm e,inter-chain}$, whose form depends on the inter-chain coupling configuration in the 2D multichains. In this paper, we consider the two simplest types of the configurations, which are named as the square type (ST) [Fig. 1(a)] and the triangle type (TT) [Fig. 1(b,c)]. The corresponding Hamiltonian can be respectively written as
$H_{\rm e,inter-chain}=$ $\zeta\sum^{L-1}_{k=1}\sum_{i=1}^{N}(\sigma_{k,i}^{+}\sigma_{k+1,i}^{-}+{\rm H.c.})$ for ST, (5a)
$H_{\rm e,inter-chain}=$ $\zeta\sum^{P}_{k=1}\sum_{i=1}^{N-1}[\sigma_{2k-1,i}^{+}(\sigma_{2k,i}^{-}+\sigma_{2k,i+1}^{-})+\sigma_{2k-1,N}^{+}\sigma_{2k,N}^{-}+{\rm H.c.}]+$
$H_{\rm e,inter-chain}=$ $\zeta\sum^{Q}_{k=1}\sum_{i=2}^{N}[\sigma_{2k,1}^{+}\sigma_{2k+1,1}^{-}+\sigma_{2k,i}^{+}(\sigma_{2k+1,i-1}^{-}+\sigma_{2k+1,i}^{-})+{\rm H.c.}]$ for TT, (5b)
where for TT the upper bounds of $L$ for odd and even term are respectively
$P=$ $(L-1)/2$ $L$ is odd, (6a)
$P=$ $L/2$ $L$ is even, (6b)
$Q=$ $(L-1)/2$ $L$ is odd, (7a)
$Q=$ $(L-2)/2$ $L$ is even, (7b)
and $\zeta$ denotes the inter-chain coupling strength, namely the interaction strength between two nearest-neighbor TLSs in two nearest-neighbor chains. Therefore, the multichains have six possible configurations based on the permutation and combination of the inter-chain and the intra-chain coupling types, which are expressed as TC ST, TC TT, SSH ST, SSH TT, TC-SSH ST, and TC-SSH TT. The TC ST, TC TT, and SSH ST are schematically depicted in Fig. 1(a), Fig. 1(b), and Fig. 1(c) respectively. Correspondingly, the SSH TT is that the TC chains in the TC TT are substituted by the SSH chains. Particularly, the TC and SSH chain are staggered arrangement in TC-SSH ST and TC-SSH TT, in which we set that the odd chains are TC type and the even chains are SSH type.
In the rotating frame of the driving frequency $\omega$, the radiation part of cavity is described by $H_{\rm c}$ as
$H_{\rm c}=-\Delta_{\rm c}c^{\dagger}c,$ (8)
where $c^{\dagger}=|c\rangle\langle 0|$ ($c=|0\rangle\langle c|$) is the photon creation (annihilation) operator, the cavity mode is decided by the frequency $\omega_{\rm c}$, $\Delta_{\rm c}=\omega-\omega_{\rm c}$ is the detuning of the cavity field frequency $\omega_{\rm c}$ with respect to the $\omega$, $|c\rangle$ is the single-photon state that only the cavity field is excited, and $|0\rangle$ is the vacuum state of the exciton-cavity system. In order to satisfy the resonant condition between the cavity and the TLSs, we set $\omega_{\rm c}=\omega_{\rm a}$, hence $\Delta_{\rm a}=\Delta_{\rm c}=\Delta$.
In the rotating frame of $\omega$ the coupling between the multi-chain aggregate and the cavity is given by
$H_{\rm ec}=\sum^{L}_{k=1}\sum_{i=1}^{N}g_{k,i}(\sigma_{k,i}^{+}c+{\rm H.c.}),$ (9)
where $g_{k,i}$ describes the coupling strength between the TLS at the $i$th site in the $k$th chain and the cavity field. For the sake of simplicity, we assume that $g_{k,i}$ is uniform for all TLSs, that is $g_{k,i}=g$. Generally, $g$ depends on the collective Rabi splitting $\Omega_{\rm R}=2\sqrt{\sum_{k=1}^{L}\sum_{i=1}^{N}g^{2}_{k,i}}=2g\sqrt{LN}$.
In order to investigate the excitation dynamics, we introduce a external coherent driving in the exciton-cavity system. In this paper we choose two driving means, i.e., the driving field is exerted on the first site of the $k$th chain as well as simultaneously on this site and the cavity, which can be described respectively as
$H_{\rm d}=$ $H_{\rm dT}=\xi_{{\rm T},k}(\sigma_{k,1}^{-}e^{i\omega t}+\sigma_{k,1}^{+}e^{-i\omega t}),$ (10a)
$H_{\rm d}=$ $H_{\rm dTc}=H_{\rm dT}+\xi_{\rm c}(ce^{i\omega t}+c^{\dagger}e^{-i\omega t}),$ (10b)
where $\omega$ is the driving frequency, $\xi_{{\rm T},k}$ ($\xi_{\rm c}$) is the strength of the TLS (cavity) drive. The second mean is denoted by ”TLS+cavity” in the following. Here, we set $k=1$, namely, only the first site in the first chain is driven. The two driving forms were used to study the two-photon blockade in the cavity QED system hamsen2017two . In the rotating frame of $\omega$, the Eq. (10) can be written as
$H_{\rm d}=$ $H_{\rm dT}=\xi_{{\rm T},1}(\sigma_{1,1}^{-}+\sigma_{1,1}^{+}),$ (11a)
$H_{\rm d}=$ $H_{\rm dTc}=H_{\rm dT}+\xi_{c}(c+c^{\dagger}).$ (11b)
When the driving field is considered, the total Hamiltonian of the open quantum system is
$H=H_{0}+H_{\rm d}.$ (12)
[…]
II.2 von Neumann entropy
In order to find the energy level crossings and anticrossings of Hamiltonian (1) in the parameter space of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$, one can calculate the von Neumann entropy of the excitons by
$S(\rho_{\rm e})=-{\rm tr}(\rho_{\rm e}{\rm log}_{2}\rho_{\rm e}).$ (13)
For the hybrid system consisting of excitons and photons, its eigenstate can be indicated as $|\psi_{\rm e-p}\rangle$. Thus the reduced density matrix $\rho_{\rm e}$ of the excitons is obtained by tracing out the degrees of freedom of the photons, i.e., $\rho_{\rm e}={\rm tr}_{\rm p}(|\psi_{\rm e-p}\rangle\langle\psi_{\rm e-p}|)$. According to the conclusion of Ref. hu2018effect , the crossings (anticrossings) correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy with parameter. Therefore, the entanglement between upper and lower energy levels is maximal at the anticrossings, at which the LZ transitions occur. The limiting of distance between two anti-cross energy levels is crossing that means the energy degeneracy. At the crossings, the excitations are exchanged coherently between the multichains and the cavity. As a result, in energetically degenerate state the exciton transport can be observed tagarelli2023electrical and boosted lin2012efficient .
II.3 Transport dynamics
Under the Born-Markov approximations and the rotating-wave approximation, the dynamics of the system that exists dissipation at each TLS in the multichains, the coupling between the first site of the first chain and the driving field as well as between the last site of each chain and the reservoir, are governed by the Lindblad quantum master equation lindblad1976generators . Then the exciton transport dynamics in the system can be investigated by the master equation
$\dot{\rho}=i[\rho,H]+\sum^{L}_{k=1}(D_{{\rm d},k}[\rho(t)]+D_{{\rm R},k}[\rho(t)]),$ (14)
where the first term in the righthand side accounts for the free evolution under the total Hamiltonian (12), the second term characterizes the decay of excitons due to the spontaneous emission with
$D_{{\rm d},k}[\rho]=\sum^{N}_{i=1}\gamma\mathcal{L}_{\sigma^{-}_{k,i}}[\rho]+\kappa\mathcal{L}_{\rm c}[\rho],$ (15)
where $\gamma$ and $\kappa$ are the decay rates of the TLSs and the cavity photons respectively, and $\mathcal{L}_{x}[\rho]=x\rho x^{\dagger}-\dfrac{1}{2}\left\{\rho,x^{\dagger}x\right\}$ is the Lindblad superoperators with standard form. The term $D_{{\rm R},k}[\rho(t)]$ in Eq. (14) represents the exciton exchange between the last site of the $k$th chain and the exciton drain, which is given by
$D_{{\rm R},k}[\rho]=\gamma_{\rm R}\bar{n}_{\rm R}\mathcal{L}_{\sigma_{k,N}^{+}}[\rho]+\gamma_{\rm R}(\bar{n}_{\rm R}+1)\mathcal{L}_{\sigma_{k,N}^{-}}[\rho],$ (16)
where $\gamma_{\rm R}$ is the exciton exchanging rate between the last site of each chain and the drain, $\bar{n}_{\rm R}$ is the average exciton number in the drain. The last site of each chain exists coupling with the drain, so $k\in\{1,2,...,L\}$. In Eq. (16), the two terms denote respectively the reflection of excitons back to the last site of the $k$th chain and the exciton that transfer from the last column TLSs of the multichains to the drain. In order to show the net transportation of exciton, we set as $\bar{n}_{\rm R}=0$ which is reasonable as the photosynthesis process in nature. Therefore, the exciton is unidirectionally transferred from the last column of the multichains to the drain without exciton back.
II.4 Efficiency of exciton transport
Figure 2: (color online) The energy spectra of the exciton-cavity coupling system as a function of the inter-chain coupling strength $\zeta$ for different configurations TC ST (a), SSH ST (b), TC-SSH ST (c), with the colormap representing the Hopfield coefficients of the content of cavity photons by solving the eigenvalues and eigenstates of the Hamiltonian (1) in zero- and single-excitation subspace. (d-f) The corresponding Hopfield coefficients of the content of cavity photons represented by line map. (g-i) The corresponding von Neumann entropy $S$ of excitons by calculating Eq. (13). The cavity-dressed energy level crossings (anticrossings) producing obvious effect on exciton transport are marked by the black (violet) vertical lines, which correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy. The corresponding parameters are $L=2$, $N=8$, $\Delta_{\rm a}=\Delta_{\rm c}=0$ eV, $h_{\rm TC}=1$ eV, $h_{\rm SSH}=1$ eV, $\delta=0.5$, and $\Omega_{R}=1$ eV.
Figure 3: (color online) Similar as Fig 2, but for different configurations TC TT (a, d, g), SSH TT (b, e, h), and TC-SSH TT (c, f, i).
Figure 4: (color online) The energy spectra of the exciton-cavity coupling system to even chain $N=8$ as a function of the dimerization parameter $\delta$ for different configurations SSH ST (a), SSH TT (b), TC-SSH ST (c), and TC-SSH TT (d). The corresponding Hopfield coefficients of photons (e-h) and von Neumann entropy of excitons (i-l). The others are similar as the Fig. 2. The parameter used is $\zeta=0.1$ eV, the others are the same as those in Fig. 2.
Figure 5: (color online) Similar as Fig 4, but for odd chain $N=9$.
To show the exciton transport dynamics of the system containing a 2D multichains and a cavity, the exciton transport efficiency is defined and derived in this subsection. Generally, the exciton transport efficiency is characterized by the ratio of the outgoing $(I^{\rm s}_{\rm o})$ and input $(I^{\rm s}_{\rm i})$ exciton current for the steady state liu2019vibration ; wei2019enhanced , that is
$\eta=\lim_{t\to\infty}\frac{I_{\rm o}}{I_{\rm i}}=\frac{I_{\rm o}^{\rm s}}{I_{\rm i}^{\rm s}},$ (17)
where the superscript ”s” refers to the steady state. On the other hand, the exciton occupation number can display the transport machanism and be used to solve the $I_{\rm i}$ and $I_{\rm o}$. The exciton occupation number on the $j$th site of the $k$th chain is given by
$p_{k,j}(t)\equiv{\rm Tr}\{\rho\sigma^{+}_{k,j}\sigma^{-}_{k,j}\}.$ (18)
Then combining Eq. (14) with Eq. (18), the equation that $p_{k,j}(t)$ obeys is derived as
$\dot{p}_{k,j}(t)=$ $i{\rm Tr}\{[\rho,H]\sigma^{+}_{k,j}\sigma^{-}_{k,j}\}+$
${\rm Tr}\{\sum_{\mu={\rm d,R}}\sum^{L}_{k=1}D_{\mu,k}[\rho]\sigma^{+}_{k,j}\sigma^{-}_{k,j}\},$ (19)
where the trace is operated overall the whole Hilbert space of the total system. The lefthand side of Eq. (19) means the rate of the change of occupation number with time, thus the first term (the second term) of the righthand side represents the input (outgoing) exciton current at the $j$th site in the $k$th chain. According to the explanation, under the TLS and TLS+cavity drive the input exciton current of the multichains can be written respectively as
$I_{\rm i}(t)=$ $i{\rm Tr}\{[\rho,H_{\rm dT}]\sigma^{+}_{1,1}\sigma^{-}_{1,1}\},$ (20a)
$I_{\rm i}(t)=$ $i{\rm Tr}\{[\rho,H_{\rm dT}]\sigma^{+}_{1,1}\sigma^{-}_{1,1}+[\rho,\xi_{c}(c+c^{\dagger})]c^{\dagger}c\},$ (20b)
which is related to the exciton injection at the first site in the $k$th chain or/and in cavity via the external coherent field per unit time. Substituting Eq. (10) for $H_{\rm dc}$ and $H_{\rm dTc}$ in Eq. (20), under the zero- and single- excitation subspace, the input exciton current is simplified to
$I_{\rm i}(t)=$ $i\xi_{{\rm T},1}(\langle 1,1|\rho|0\rangle-{\rm H.c.}),$ (21a)
$I_{\rm i}(t)=$ $i\xi_{{\rm T},1}(\langle 1,1|\rho|0\rangle-{\rm H.c.})+i\xi_{\rm c}(\langle c|\rho|0\rangle-{\rm H.c.}).$ (21b)
According to the Hermitian property of density matrix, under the TLS and TLS+cavity drive the input exciton current is finally written as
$I_{\rm i}(t)=$ $2\xi_{{\rm T},1}{\rm Im}[\langle 0|\rho|1,1\rangle],$ (22a)
$I_{\rm i}(t)=$ $2\xi_{{\rm T},1}{\rm Im}[\langle 0|\rho|1,1\rangle]+2\xi_{c}{\rm Im}[\langle 0|\rho|c\rangle].$ (22b)
Analogous to the input exciton current, the outgoing exciton current can be defined as the outflow of exciton transfering from the last column in the multichains to the drain per unit time. According to the means of the third term of righthand in Eq. (19), it can be given by
$I_{\rm o}(t)={\rm Tr}[\sum^{L}_{k=1}D_{{\rm R},k}[\rho(t)]\sigma_{k,N}^{+}\sigma_{k,N}^{-}],$ (23)
where each chain has exciton outgoing to the drain for all configurations, thus $k\in\{1,2,...,L\}$. Pluging the Eq. (16) into the Eq. (23), under the zero- and single- excitation assumption, the form of the outgoing exciton current is simplified as
$I_{\rm o}(t)=\gamma_{\rm R}\sum^{L}_{k=1}p_{k,N},$ (24)
where $p_{k,N}$ is the occupation number at the last site in the $k$th chain.
III Spectrum of coupling system
There are a large number of studies on the spectra of the TC and SSH Hamiltonian for the open boundary conditions campos2021generalization ; sirker2014boundary ; nie2020bandgap ; wei2022cavity ; cai2019observation . The TC chain can be regarded as a special configuration of the SSH chain. For the SSH chain, the topologically trivial and nontrivial phases depend on the parity of the number of an 1D TLSs chain and the qubit-qubit coupling configuration. According to the Ref. cai2019observation , the dimerization parameter $\delta>0$ ($\delta<0$) corresponds to the topological winding number $\nu=1$ ($\nu=0$), i.e., the topologically nontrivial (trivial) regime. In Ref. sirker2014boundary , the results show that there is one zero-energy mode in whole range of $\delta$ and are two localized nearly-zero-energy modes in the topologically nontrivial regime $\delta>0$ for the spectra of odd and even chain respectively. The finite gap between the two edge modes leads to the formation of the hybrid edge modes that overlap with each other, generating the polaritonic states when the edge modes are dressed by cavity due to the coupling between the excitons and the photons. Comparing to the single chain system, the presence of the inter-chain coupling induces the extension of the spectra of the 2D multi-chain system. As we all known, the gap corresponds to the energy level anticrossings, at which the LZ transitions among the levels occur in multi-level LZ system, adding greatly to the complexity of the system dynamics hu2018effect . What’s more, the energy level crossings mean that the corresponding quantum states are degeneracy. In the exciton-cavity coupling system, at the crossings the excitations are coherently exchanged between the TLSs and the cavity. Therefore, the crossings and anticrossings of the edge modes dressed by cavity can lead to the abundant exchange of photon between the multichains and the cavity, increasing or decreasing polaritons. The forming polaritons act as intermediary in exciton transport, where the excitons are transported among site $(k,1)~{}\rightarrow$ cavity $\rightarrow$ site $(k,N)$ without experiencing the dissipation of the other TLSs on multichains wei2022cavity . Therefore, the generating of the polaritons leads to the exciton transport efficiency and currents increasing, which can be controlled by the crossings and anticrossings.
Figure 6: (color online) The dependence of the steady-state occupation numbers (a1 and b1) $p_{11}$ on site 11, (a2 and b2) $p_{\rm bulk}$ on bulk sites, (a3 and b3) input exciton current $I_{\rm i}$, (a4 and b4) outgoing exciton current $I_{\rm o}$, and (a5 and b5) transport efficiency $\eta$ on the inter-chain coupling strength $\zeta$ (a1-5) and the dimerization parameter $\delta$ (b1-5) for different configurations without cavity. In (b1-5) the solid and dash line correspond respectively to $N=8$ and 9. The parameters used are $\zeta=0.1$ eV, $\xi_{{\rm T},1}=\xi_{\rm c}=0.1$ eV, $\gamma_{\rm d}=0.01$ eV, $\gamma_{\rm R}=0.1$ eV, and the others are the same as those in Fig. 2.
Figure 7: (color online) The dependence of the steady-state dynamics of the hybrid system on the inter-chain coupling strength $\zeta$ in the presence of cavity under TLS drive (a1-6) and TLS+cavity drive (a1-6). (a5-b5) The occupation number of photon with $\zeta$, in which the inflection points are marked by the vertical fine lines, and the others are similar to Fig. 6. The parameter used is $\kappa=0.01$ eV, and the others are the same as those in Fig. 6.
Figure 8: (color online) The dependence of the steady-state dynamics of the hybrid system on the dimerization parameter $\delta$, and the others are similar to Fig. 7. The parameters used are the same as those in Fig. 7.
In order to analyse the spectra of the coupling system consisting of the multichains and the cavity, its eigenvalues can be obtained by numerically diagonalizing the Hamiltonian (1). On the other hand, the formation of polaritons can be investigated by calculating the Hopfield coefficients $|\alpha_{k,i;c}|^{2}$. When the $j$th normalized eigenvector is expanded as $|\psi_{j}\rangle=\Sigma_{k}^{L}\Sigma_{i}^{N}\alpha_{k,i}|k,i\rangle+\alpha_{c}|c\rangle$, $|\alpha_{k,i;c}|^{2}$ denotes Hopfield coefficients, which characterize the corresponding hybridizing weight of different excitonic and photonic contents. Hence, the color maps Fig. 2 (a-c), 3 (a-c) and Fig. 4 (a-d), 5 (a-d) show respectively the spectra of the exciton-cavity coupling system as functions of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$ for different coupling configurations, containing the photonic Hopfield coefficients $|\alpha_{c}|^{2}$. To show obviously the photonic Hopfield coefficients, we plot the line maps of $|\alpha_{c}|^{2}$ of each eigenstate as functions of $\zeta$ and $\delta$ in Fig. 2 (d-f), 3 (d-f) and Fig. 4 (e-i), 5 (e-i) respectively. To find out the crossings and anticrossings in spectra, the corresponding von Neumann entropy $S$ of excitons is calculated through Eq. (13), and is shown in Fig. 2 (g-i), 3 (g-i), 4 (i-l), and 5 (i-l). One can see that the zero-energy modes exist for each situation. It is obvious that only the nearly-zero-energy modes are distinctly dressed by the cavity mode, generating corresponding polaritonic branches, because the high excited states are hardly dressed for the given coupling strength between excitons and photons. At the crossings (anticrossings) one can see that the photonic contents exchange between the upper and lower polaritonic branches with $\zeta$ and $\delta$, but the distributed ratio between the upper and lower polaritonic branches depends on the locations of the crossings (anticrossings). It means that the exchange of excitation or the LZ transition of photon occurs, inducing the change of the exciton transport efficiency and currents, which will be illustrated in the next section by the steady-state dynamics. Further, it can be discovered that the effect of the inter-chain coupling strength on the crossings (anticrossings) is more significant than that of the dimerization parameter. The number of crossings (anticrossings) affecting transport is relate to the coupling configuration in the range of $\zeta$. Comparing to the changes of $|\alpha_{c}|^{2}$ and of $S$ with $\zeta$ and $\delta$, a common phenomenon can be seen that the locations of the intersections of the photonic Hopfield coefficients and of the von Neumann entropies corresponding to different eigenstates are consistent.
In the next section, the steady-state dynamics with $\zeta$ and $\delta$ will be shown, in which the relation between the dressed crossings (anticrossings) and the occupation number of photons can be analysed. The changes of exciton transport currents and efficiency can be explained. We aim at revealing the mechanism of exciton transport by finding the optimal range of parameters on exciton transport and comparing the spectra and steady-state dynamics for different configurations.
Figure 9: (color online) The dependence of the steady-state dynamics on the inter-chain coupling strength $\zeta$ and the detuning $\Delta$ of the driving frequency to the TLS without cavity under TLS drive (a-f, ab-fb), and with cavity under TLS drive (g-l, gc-lc) and TLS+cavity (the TLS and cavity drive exist simultaneously) drive (m-r) for different configurations. The corresponding physical quantities of each row are shown at right. The zones that cavity produces significant effect on exciton transport efficiency are marked by ellipses in (g-l), the corresponding zone are marked in (ab-fb) and (gc-lc). The parameters used are the same as those in Fig. 7.
Figure 10: (color online) The dependence of the steady-state transport efficiency $\eta$ on the dimerization parameter $\delta$ and the detuning $\Delta$ of the driving frequency relative to the TLS, in which the corresponding conditions are written at the top and the right of this figure. The parameters used are the same as those in Fig. 7.
Figure 11: (color online) The steady-state dynamics as a function of the number $L$ of the chains without (a1111-g1111) and with (a2222-h2222) cavity for different configurations, in which $p_{1,i}$, $p_{L,i}$, and $p_{bulk-in}$ are respectively represent the occupation number of the first row, the last row, and the bulk state denoted as ”bulk-in” containing all TLSs inside of multi-chain system. The parameter used is $N=10$, and the others are the same as those in Fig. 7.
Figure 12: (color online) The steady-state dynamics as a function of the length $N$ of chain for different configurations and number of the chains, (a-d) without cavity under TLS drive, (e-h) with cavity under TLS drive. The 1,2,3,4th column correspond to $L=2,3,4,5$ respectively. The parameters used are the same as those in Fig. 7.
IV The exciton transport dynamics by multichains
In this section, we discuss the effects of the inter-chain coupling, the dimerization parameter, the number and length of the chain on the exciton transport dynamics by numerically solving the master equation Eq. (14) under an external driving, which are displayed in two subsections below. According to the analysis in the Sec. III, the nearly-zero-energy modes are dressed by cavity for strong exciton-cavity coupling, namely the polaritonic branches forming. In order to let the polaritons act as a dominated role in exciton transport, we set that the Rabi splitting is far stronger than the decay rates of the cavity and the TLSs, i.e., $\Omega_{R,k}\gg(\kappa,\gamma)$.
IV.1 The effects of the inter-chain coupling and the dimerization parameter
Here, we investigate primarily the corresponding relation between the energy level crossings (anticrossings) dressed by cavity and exciton transport dynamics to explain the exciton transport mechanism. Based on the relation and the properties of exciton transport dynamics, the optimal parameter ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$, as well as the optimal coupling configurations are expected to be presented.
Fig. 6 shows the dependence of the steady-state exciton transport dynamics of the multi-chain system without cavity on $\zeta$ and $\delta$ for the six chosen configurations. From the Eq. (24), it can be known that the outgoing exciton current $I_{\rm o}$ is directly proportional to the total occupation number $\Sigma_{1}^{L}p_{k,N}$ of the outgoing sites, thus only $I_{\rm o}$ is shown in following figures. Hereinafter $\Sigma_{1}^{L}p_{k,N}$ is denoted by $p_{N}$. One can see that the effects of the inter-chain coupling and the dimerization parameter on exciton transport are significantly different, in which the steady-state occupation number $p_{11}$ of site ”11” (The two ”1” denote respectively the 1st row and the 1st column in the array of the TLSs shown in Fig. 1. This representation is similar below, unless otherwise noted.), $p_{\rm bulk}$ of bulk sites (The ”bulk” denotes the all sites from the 2nd to the $(N-1)$th column), the input and outgoing exciton current $I_{\rm i,o}$, and the exciton transport efficiency $\eta$ with $\zeta$ are symmetric about $\zeta=0$ in $[-1,1]$, but with $\delta$ are not.
For the six configurations, the descending sort of efficiency is successively TC-SSH ST, TC ST, TC TT, TC-SSH TT, SSH ST, SSH TT in $[-1,1]$ [Fig. 6(a5)]. For the TC-SSH ST, near $\zeta=0.3$ the sites of bulk [Fig. 6(a2)], input, and outgoing possess commonly relatively high occupation numbers of exciton, which support exciton transport on multichains. Therefore, the exciton transport efficiency, input and outgoing currents are the largest [Fig. 6(a3,4,5)]. On the contrary, for $\zeta=0$ and $\zeta\rightarrow\pm 1$, although the site 11 possesses more excitations [Fig. 6(a1)], the exciton occupation number $p_{N}$ at outgoing sites is too few to transfer to the drain, leading the low efficiency and exciton currents. Specially, the excitons are distributed mainly at bulk sites for $\zeta\rightarrow\pm 1$ [Fig. 6(a2)], which is beneficial to energy storage and can be used to design quantum battery quach2022superabsorption . Except for the SSH ST, the other configurations also have the property for $\zeta\rightarrow\pm 1$, as well as TC TT and TC-SSH TT also do near $\zeta=\pm 0.4$. However, for TC ST near $\zeta=\pm 0.35$ and ±$\pm 0.97$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ are relatively high, thus the ability both of energy transport and storage are strong. Near $\zeta=\pm 0.6$, although the exciton transport efficiency is large, exciton currents are relatively small. Thus it is not beneficial to the exciton transport.
It has four configurations SSH ST, SSH TT, TC-SSH ST, TC-SSH TT containing SSH chain. Because of the different spectra between odd and even SSH chain, in Fig. 6(b1-5) there are two situations $N=8,~{}9$ about the steady-state dynamics with $\delta$ in the absence of cavity. It is obvious that the exciton currents and transport efficiency are the largest near $\delta=0.3$ for the configurations SSH ST and TC-SSH ST of $N=8$ than that in the other range of $\delta$ and for the other configurations [Fig. 6(b3-5)], which are good for effective exciton transport. However, the SSH TT of $N=9$ is the most beneficial to energy storage because of the high occupation number $p_{\rm bulk}$ of bulk sites [Fig. 6(b2)] as well as the lowest exciton currents and transport efficiency. As shown in Fig. 6(b5), the exciton transport efficiency is high for $\delta<0$, but the exciton currents approach to zero, thus there is no real exciton transport. The reason is that the site 11 can not be coherently excited by drive, namely the excitons can not be injected in multichains [Fig. 6(b1,2)]. When $\delta>0.35$, the exciton currents and transport efficiency decrease sharply with $\delta$ or too small, because the excitons are localized at the left-edge site so that the occupation number of the right-edge sites is small for the TC-SSH ST and TC-SSH TT of $N=8,9$ or approaches to zero for the others.
In order to investigate the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ on the exciton transport dynamics in the presence of cavity, we plot the Fig. 7 and 8 corresponding to Fig. 6, in which the occupation number $p_{\rm c}$ of photons is added. We pay closely attention to the change of the steady-state dynamics with $\zeta$ and $\delta$ at the crossings (anticrossings) of nearly-zero-energy modes dressed by cavity.
Under the TLS and the TLS+cavity drive, the effect of the inter-chain coupling $\zeta$ on the steady-state dynamics of the multi-chain system with cavity is presented for different configurations in Fig. 7(a1-6) and 7(b1-6) respectively. Particularly, to analyse what role the cavity acts as in exciton transport, we plot the dependence of the occupation number $p_{\rm c}$ of photons on $\zeta$ and mark its inflection points by the vertical fine lines in Fig. 7(a5), which correspond to the locations marked by the vertical fine lines in Fig. 2 and 3. Comparing Fig. 2, 3 with Fig. 7(a5), one can discover that each inflection point has corresponding crossings or anticrossings with the same locations. Next, we mainly analyse the effect of cavity on the exciton currents $I_{\rm i,o}$ and transport efficiency $\eta$. It can be seen that the presence of cavity breaks the symmetry of $I_{\rm i,o}$ with respect to $\zeta=0$ for configurations TC ST, SSH ST, SSH TT, and TC-SSH ST [Fig. 7(a3,4)]. Comparing to Fig. 6, it can be seen that the exciton transport is suppressed by cavity at the range of $\zeta\neq 0$. For example, near $\zeta=-0.35$ $p_{\rm c}$ is high [Fig. 7(a5)], suppressing the exciton currents in the main channel with the multichains for TC ST [Fig. 7(a3,4)]. On the other hand, there are three configurations SSH ST, SSH TT, and TC-SSH TT for which the cavity induces significant increase of $I_{\rm i,o}$ and $\eta$ near $\zeta=0$. In the situation, the exciton transport is assisted by polaritons, in which $p_{\rm c}$ is high, leading the increase of exciton occupation number on the last column sites $p_{N}$ [Fig. 7(a4)]. In fact, $\zeta=0$ means that only one chain undertakes the exciton transport, in which polariton-assisted exciton transport is demonstrated wei2022cavity . Especially, it can be seen obviously that the peaks of $p_{\rm c}$ and $I_{\rm i,o}$ are at different locations. The reason is that at the peak of $p_{\rm c}$ there is not enough excitations to be injected in multichains, causing relatively low $I_{\rm i,o}$. Because of few excitations injected [Fig. 7(a1)] for TC ST, TC TT and few $p_{N}$ [Fig. 7(a4)] for TC-SSH ST at the locations of the peaks of $p_{\rm c}$ cavity has no effect on exciton transport near $\zeta=0$. As shown in Fig. 7(b5), adding the cavity drive $p_{\rm c}$ increase largely, but $I_{\rm i,o}$ and $\eta$ do not. Near $\zeta=0$, because of the transfer of excitations the occupation number $p_{11}$ of site 11 increases [Fig. 7(b1)], causing the decrease of $p_{\rm c}$ for SSH ST, SSH TT, TC-SSH ST, and TC-SSH TT [Fig. 7(b5)], suppressing the exciton transport of the main channel with porlaritons [Fig. 7(b3,4,6)]. However, one can see by comparing Fig. 7(b3,4) with Fig. 7(a3,4) that the increase of $p_{\rm c}$ suppresses significantly the exciton currents of the main channel with the multichains near $\zeta=0.2$ for TC-SSH ST and near $\zeta=0.35$ for TC ST. In a word, the competition between cavity photons and excitons is the key affecting exciton transport.
Fig. 8(a1-6) and 8(b1-6) show the dependence of the steady-state dynamics on the dimerization parameter $\delta$ under the TLS and the TLS+cavity drive in the presence of cavity respectively. We analyse firstly the effects of $\delta$ and cavity on exciton transport under the TLS drive. Although the exciton transport efficiencies are centered approximatively on $\delta=0$ [Fig. 8(a6)], there are no the exciton currents in $\delta<0$ except for the configurations SSH ST and TC-SSH ST of $N=9$ [Fig. 8(a3,4)]. The root cause is that excitons can not enter to multichains through the site 11 [Fig. 8(a1,3)]. It can be seen that for the configurations SSH ST and TC-SSH ST of $N=8$ in the range $0.350<\delta<$ 0.35 the presence of cavity suppresses significantly the exciton currents $I_{\rm i,o}$ [Fig. 8(a3,4)] and transport efficiency $\eta$ [Fig. 8 (a6)] by comparing to Fig. 6(b3-5). Particularly, at $\delta=0.26$ for SSH ST and $\delta=0.36$ for TC-SSH ST of $N=8$, the troughs of $I_{\rm i,o}$ and $\eta$ appear, which correspond to the peak and the trough of $p_{\rm c}$ [Fig. 8(a5)] and to the crossings near zero-energy modes dressed [Fig. 4] marked by the vertical fine lines, respectively. The reason for the former is that the excitation on the first and the last column TLSs is transferred to cavity through coherent exciting at cavity dressed crossings as shown in Fig. 8(a1,3,4), suppressing the exciton currents and transport efficiency on the main channel with multichains. But for the latter, the occupation number $p_{N}$ of exciton on the last column TLSs is too few, limiting the exciton transport. At $\delta>0.26$ for SSH ST and $\delta>0.36$ for TC-SSH ST of $N=8$, the limited photon number $p_{\rm c}$ [Fig. 8(a5)] and the decreased occupation number $p_{\rm bulk}$ on bulk sites [Fig. 8 (a2)] lead to that the polariton becomes the main channel, thus the exciton transport is strengthened by cavity comparing with Fig. 6(b3,4). However, for TC-SSH TT both the multichains and the polariton dominate the exciton transport because of the high $p_{\rm bulk}$ and increasing $p_{\rm c}$ with $\delta$. For the odd chains, in about $\delta>0.35$ the low $p_{N}$ is main reason limiting exciton transport. When the cavity drive is added based on the TLS drive, there are distinct effects on exciton transport in $0.350<\delta<$ 0.35 for SSH ST and TC-SSH ST of $N=8$. The increased $p_{\rm c}$ induces the increase of polaritons [Fig. 8(b5)], enhancing the exciton transport [Fig. 8(b3,4)] than that under TLS drive [Fig. 8(a3,4)], but in $\delta>0.35$ because of the low $p_{\rm c}$ and $p_{\rm bulk}$ [Fig. 8(b2)] the exciton transport is limited for SSH ST. Interestingly, the exciton currents and transport efficiency are significantly large in $\delta<0$ for SSH ST and TC-SSH ST of $N=9$, which break the universal situations that the exciton currents and transport efficiency are higher in $\delta>0$ than that in $\delta<0$. Specially, when the cavity is driven, the occupation number of site 11 is not necessary to transport exciton efficiently [Fig. 8(b1)]. In the condition, the high $p_{\rm bulk}$ [Fig. 8(b2)] and $p_{\rm c}$ [Fig. 8(b5)] illustrate that both the multichains and the polaritons participate the exciton transport. Although $p_{11}$ no longer is the condition restricting exciton transport under the TLS+cavity drive, $p_{N}$ still is. Therefore, the low $p_{N}$ suppresses the exciton transport for SSH TT and TC-SSH TT of $N=9$ in $\delta<0$.
When the TLS transition frequency detuning $\Delta$ is considered, we plot the steady-state dynamics of the hybrid system in Fig. 9 and 10, which show respectively the effects of inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ for different configurations. Comparing with Fig. 2 and 3, the structures of exciton transport efficiency are similar to those of spectra. In Fig. 9 the zones of $\zeta$ and $\Delta$ that cavity has obviously effect on exciton transport are marked in Fig. 9(ab-fb, g-l, and gc-lc) by ellipses through comparing to Fig. 9(a-f). According to the effect these configurations can be classified into two categories. One is TC ST and TC TT that the exciton transport is suppressed, the other is the remaining configurations that the exciton transport is enhanced. For the former the transfer of excitation to cavity leads to the exciton transport channel with the multichains suppressed [Fig. 9(gc,hc)]. For the letter, as shown in Fig. 9(cb-fb) the occupation number $p_{\rm bulk}$ of bulk state is too small in the zone to support exciton transport of the channel with the multichains, thus the polaritons dominate the transport of exciton, in which the occupation number $p_{\rm c}$ of cavity photons is relatively large [Fig. 9(ic-lc)]. Further, we can illustrate the mechanism through spectra. Interestingly, the structure of $p_{\rm c}$ [Fig. 9(gc-lc)] is similar to that of the cavity-dressed energy levels [Fig. 2 and 3]. In the intensified zone by cavity there are the cavity-dressed anticrossings [Fig. 2 and 3], thus LZ transition is the foundation that cavity effects exciton transport. Particularly, in the high efficient zone of exciton transport there are two anti-crossed zonate zones with large $p_{\rm c}$ [Fig. 9(ic)] corresponding to the dressed anticrossings for SSH ST. When the cavity drive is exerted further, for the two categories of configurations the exciton transport is further suppressed and enhanced by cavity respectively. For SSH ST the effect of cavity drive on exciton transport is the biggest, in which the high efficient zone is significantly extended. In parameter space of $\delta$ and $\Delta$, as shown in Fig. 10 the steady-state dynamics depend on the coupling configuration, the cavity, and the method of drive. For SSH ST, TC-SSH ST of $N=8,~{}9$ and SSH TT of $N=8$, the cavity has apparent effects on exciton transport efficiency in the zone of the fully dimerized limit, namely, $\delta\rightarrow\pm 1$, and near the zero detuning [Fig. 10(a-d, g-i, m-p)]. Nevertheless, because the TC chain and the diagonal inter-chain coupling suppress the dimerized effect, leading to effect of cavity suppressed for TC-SSH TT [Fig. 10(s-v)]. Undoubtedly, in the zone enhancing exciton transport the polaritons act as prime role whose mechanism is coherent excitation (LZ transition) located at the cavity-dressed crossings (anticrossings) according to the analysis on Fig. 4, 5 and 8. When cavity drive is added, it can be seen that the high efficient zones are extended and the exciton transport efficiencies are strengthened in the zone of the fully dimerized limit and near the zero detuning for SSH ST, SSH TT, and TC-SSH ST of $N=8,~{}9$ [Fig. 10(e, f, k, l, q, r)].
IV.2 The effects of the number and the length of chains
In this part, we discuss the dependence of exciton transport dynamics on the number $L$ and the length $N$ of the chains, and contribute to reveal the inter-chain joint effect, to find optimal $L$, $N$, and configurations with high transport efficiency and currents in Fig. 11 and 12. It can be seen that for the three configurations SSH ST, TC-SSH ST, and TC-SSH TT the number $L$ of the chains has obvious effect in $L\leq 6$ [Fig. 11(e, f, h)]. Although there is large input exciton current, the occupation number $p_{k,N}$ of the last column is so few that the outgoing exciton current is little for SSH TT and the excitons are aggregated on the inside of the multichains with $L$ for SSH ST [Fig. 11(c1, e1, g1, d1)]. Therefore SSH ST is perfect candidate applied to quantum battery in the six configurations. What’s more, for SSH TT the most of excitons is occupied by the upper edge sites, which is similar to skin effect. In fact, excitons is restricted on the first chain with $L$ because of the low efficient transport of excitons among chains. On the contrary, because of the expand of excitons with $L$ the inner occupation number $p_{\rm bulk-in}$ increase and the occupation number of the first chain decrease for SSH ST. In the presence of cavity the steady-state dynamics display abundant phenomena as shown in Fig. 11(a2-h2). The prominent one is that the exciton currents and transport efficiency present distinct oscillation with $L$ for SSH TT and TC-SSH TT, in which $I_{\rm i,o}$ and $\eta$ are peak (trough) for odd (even) number of $L$ and $N=10$ [Fig. 11(c2, g2, d2)]. The difference between steady-state dynamics of odd and even number become small with $L$. In addition to, the exciton transport is strengthened by cavity for SSH ST.
There is an obvious character that the exciton transport efficiency $\eta$ behaves oscillation with the length $N$ of the chains in Fig. 12. We firstly discuss the situation without cavity in Fig. 12(a-d). The oscillation is relative weak for TC ST and depends on the parity of $L$ and $N$ for TC TT in the absence of cavity. Thus the diagonal inter-chain coupling induces difference between exciton transports of odd and even $N$ in long TC TT multichains as shown in Fig. 12(b,d). It can be found that because of the dimerization the prominent oscillation with $N$ exists in the short multichains, but disappears in the long multichains for SSH ST and SSH TT [Fig. 12(a-d)]. In fact, the dimerization hinders the exciton diffusing for later. However, in the mixed multichains of TC-SSH because of the presence of TC chains the dimerization is suppressed, the oscillation is enhanced for TC-SSH ST and TC-SSH TT. In the two configurations the oscillation is consistent for odd $L$ and is opposite for even $L$ each other about the parity of $N$, such as $\eta$ of $N=16$ is trough and that of $N=17$ is peak for TC-SSH TT when $L=2$, but it is inverse for TC-SSH ST. When the cavity is considered the oscillation is significantly strengthened for the configurations containing SSH chain, as shown in Fig. 12(e-h). For SSH ST the oscillation decreases with $L$, which is caused by the dissipation of the more and more TLSs. Whatever situation the oscillation of $\eta$ with $N$ for SSH ST and SSH TT is consistent. Interestingly, it behaves inverse dependence on the parity of $L$ between SSH TT and TC-SSH ST, TC-SSH TT, i.e. for the former the oscillation amplitude of odd $L$ is bigger than that of even $L$, and for the latter the oscillation amplitude of even $L$ is bigger than that of odd $L$. Furthermore, the dependence of the exciton transport efficiency on the parity of $N$ can be changed by $L$ and cavity for TC-SSH ST and TC-SSH TT. For example, comparing with $L=2$ and $L=3$ for TC-SSH ST $N=16$, the $\eta$ corresponds respectively to peak and trough [Fig. 12(a, b)], as well as comparing with Fig. 12(b) and (f) for TC-SSH ST $N=14$ it corresponds respectively to trough and peak. We also note by comparing with Fig. 12(b) and (f), as well as $L=3$ [Fig. 12(f)] and $L=4$ [Fig. 12(g)] for TC-SSH ST, that the period of oscillation with $N$ also is changed by cavity and $L$.
V Summary
In summary, the steady-state dynamics of exciton in the hybrid system consist of a 2D multichains and a cavity are studied through the spectra and the Lindblad quantum master equation, in which six inter-TLS coupling configurations are chosen as TC ST, TC TT, SSH ST, SSH TT, TC-SSH ST and TC-SSH TT based on the TC and SSH model, as well as square and triangle type. The effects of the inter-chain coupling strength, the dimerization parameter, the number of chain, and the length of chain on the spectra and the steady-state dynamics are investigated. The results show that in the absence of cavity the exciton transport efficiency is decided by the distribution of population of exciton on the whole multichains. Besides, when the cavity is considered the occupation number of the cavity photons depend significantly on the cavity-dressed energy level crossings and anticrossings near the zero-energy modes, at which the coherent excitation and LZ transition between the multichains and the cavity occur. Therefore, the exciton transport efficiency and currents via the polariton channel are changed at the locations of the crossings and anticrossings. We can conclude that the exciton transport dynamics are controlled by the crossings and anticrossings. The spectra display that the locations of the crossings and anticrossings are related closely to the coupling configurations, the inter-chain coupling strength, and the dimerization parameter, as well as the stead-state dynamics also are correspondingly. Particularly, the exciton transport efficiency and currents are symmetric about the inter-chain coupling strength $\zeta=0$ in the multi-chain system without cavity, but the symmetry can be broken by cavity. In the fully dimerized limit, the presence of cavity enhances exciton transport for SSH ST, TC-SSH ST and TC-SSH TT of the even length of the chains. Interestingly, when the cavity drive is simultaneously exerted, the exciton transport is not limited by the occupation number of site $11$ and is strengthened for SSH ST and TC-SSH ST of the odd length of the chains in the negative range of the dimerization parameter. Further, in the space of the TLS transition frequency detuning and the inter-chain coupling strength we discover that the exciton transport is suppressed for TC ST and TC TT and strengthened for the other configurations by cavity at the range corresponding to the crossings and anticrossings, in which the polaritons act as vital role and the effects are enhanced by the cavity drive. In the space of the TLS transition frequency detuning and the dimerization parameter, the steady-state dynamics depend on the coupling configuration, driven method, and cavity, especially in the zone of the fully dimerized limit. Finally, it can be found that the exciton transport is obviously related to the parity of the number and length of the chains with cavity, in which the exciton transport efficiency behaves oscillation with them. The discrepancy between of the efficiency and currents for odd and even is related to the configuration. In a word, our results reveal the exciton transport mechanism of the 2D hybrid system, which provides theoretical foundation for designing controllable and high efficient exciton transport devices, such as solar cell. Of course, it has many points of view to reveal the mechanism, like topological phase transition of the system.
VI Acknowledgements
This work is supported by the Academic Ability Promotion Foundation for Young Scholars of Northwest Normal University in China under the Grant No. NWNU-LKQN2024-10, National Natural Science Foundation of China Regional Fund under Grant No. 12164042, National Natural Science Foundation of China under Grant No. 12104374, National Natural Science Foundation of China under Grant No. 12065022.