### v3

# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

Abstract
The accomplished functions of a variety of quantum devices are closely associated with the controlling of exciton transport. To this end we study the exciton transport of the two-dimensional system consisting of two-level multichains with various coupling configurations in a cavity. Two types of the chains, the homogeneous and heterogeneous coupling chain, as well as two inter-chain coupling conformations, the square and triangle type, are considered. The effects of the inter-chain coupling, the dimerization parameter, the cavity, the length and number of the chains on exciton transport are in detail investigated for different coupling configurations through the spectra, the Hopfield coefficients, and the steady-state dynamics of the system. The results show that in the absence of cavity the exciton transport currents and efficiency are decided by the distribution of exciton on the multichains. However, when the cavity is considered the exciton transport can be significantly enhanced or suppressed by the polariton at the cavity-dressed energy level crossings and anticrossings near zero-energy modes, at which the coherent excitation and Landau-Zener transitions occur. Meanwhile, we discover that the incontinuity and extreme points of the second order partial derivative of the Hopfield coefficients of the photon with the inter-chain coupling and the dimerization parameter correspond respectively to the crossings and anticrossings at the extreme points of the photon occupation number. In addition, the exciton transport efficiency is closely related to the parity of both the length and number of the chains, which behaves as oscillation. This work is important for understanding the exciton transport mechanism in the multichain-cavity system, and provides theoretical basis for excitonic devices with controllable and efficient exciton transport.
I Introduction
In quantum device, the carriers of energy and information transport are mainly electron, photon, and exciton. Enhancing and controlling the efficiency, the speed, the distance, and the current of carrier transport are the center to accomplish the functions of various quantum devices, such as organic solar cells wurfel2015impact ; lu2022simultaneously ; melianas2017charge , organic light-emitting diodes shan2011transport ; tanase2003unification ; shen2022through , quantum dot solids zhang2022ultrafast , organic semiconductor devices sneyd2021efficient ; hagenmuller2017cavity , quantum simulation, and quantum computation georgescu2014quantum ; yang2019quantum etc. The factors affecting exciton transport are various, including the dissipation cainelli2021exciton ; zhu2024quantum ; bose2022effect , exciton-exciton annihilation wittmann2020energy , disorder zhu2022single ; quenzel2022plasmon ; schachenmayer2015cavity ; allard2022disorder ; balasubrahmaniyam2023enhanced ; ribeiro2022multimode , defect sites wagner2011bound ; akselrod2014visualization , and topological property buendia2024long ; allard2023multiple ; nie2021topology . Thus, a large of studies are devoted to revealing the exciton transport mechanisms and properties of various quantum system engel2007evidence ; menke2013tailored , and to achieving the requirements for the property of exciton transport through reduce and avoid detrimental factors.
In experiment, the one-dimensional (1D) system carrying electron, photon, or exciton transport can be realized by various platforms, including organic aggregates molecule chainssu1979solitons , superconducting qubit circuits roushan2017spectroscopic ; xu2018emulating ; blais2003tunable , Rydberg-atom chains barredo2015coherent ; orioli2018relaxation , plasmons su2021perovskite ; seitz2020exciton , semiconductorssneyd2021efficient ; giannini2022exciton ; muller2023directed . At room temperature the strong exciton-photon coupling between organic aggregates and microcavities lidzey1998strong ; lidzey1999room ; kena2008strong ; tischler2005strong ; coles2014polariton or plasmons bellessa2014strong ; bellessa2004strong ; zengin2015realizing ; li2018plasmon can result in new applications like polariton lasing kena2010room and the polaritons forming between the edge state excitons and the cavity photons kavokin2003cavity ; agranovich2003cavity , in which the molecular energy structures are modified galego2015cavity . The presence of polaritons can adjust the efficiency wei2019enhanced ; wei2022cavity ; feist2015extraordinary , rate wang2021polariton ; aroeira2023theoretical , and distance du2018theory ; myers2018polariton ; sandik2024cavity ; takazawa2010fraction of the exciton transport because the decay of excitons on chains is suppressed.
The hybrid system consisting of an 1D molecule chain and a cavity can be described by the Tavis-Cummings model tavis1968exact . The chain with the staggered hopping interaction is depicted by the Su-Schrieffer-Heeger (SSH) model su1979solitons . In the SSH chain there are topologically trivial and nontrivial phases for the even number of the two-level systems (TLSs) cai2019observation . The even chain in the topologically nontrivial regime has a small gap bewteen two localized nearly-zero-energy modes in the middle of the topological bandgap, which leads to the formation of the hybrid edge modes with the left and the right edge state. Therefore, the nontrivial edge states topologically protected can enhance the exciton transport for a finite-size SSH chain, which was verified theoretically and experimentally chang2020probing ; wei2022cavity . In fact, the nature of the gap is the energy level anticrossing, at which the Landau–Zener (LZ) transition occurs. In the two-dimensional (2D) and three-dimensional tungsten disulfide material, the anti-crossing dispersions were presented, in which the boosted transport capability of excitons was corroborated in experiment guo2022boosting ; yuan2017exciton . The properties of the multichains system coupling to a cavity were theoretically investigated kvande2023finite . Thus, the potential properties of the more than 1D system on the exciton transport are fascinating and valuable to study.
Motivated by the works on the high dimensional tungsten disulfide, in this paper we devote to the studies of the inter-chain joint effect on exciton transport in 2D TLSs array with a cavity. Each chain contains two alternative coupling types, i.e., the homogeneous (HO) and the heterogeneous (HE) coupling, and two conformations of the inter-chain coupling are set, i.e., square and triangle types. Firstly, the energy spectra of the hybrid system in the condition of the strong exciton-cavity coupling with the inter-chain coupling strength and the dimerization parameter are presented by diagonalizing numerically the Hamiltonian for six configurations. Meanwhile, we calculate the von Neumann entropy of exciton to confirm the locations of the crossings and anticrossings, and the second order partial derivative of the photonic Hopfield coefficients to show the change of the occupation number of the cavity photon. What’s more, the effects of the inter-chain coupling, the dimerization parameter, the cavity, the TLS transition frequency detuning, the number and length of the chains on the steady-state dynamics of exciton transport are investigated for different configurations by numerically solving the Lindblad quantum master equation under the TLS and TLS+cavity drive. In the absence of cavity, the channel of exciton transport only is the multichains, in which its current and efficiency depend on the distribution of exciton on the whole multichains. Importantly, for the hybrid system the exciton transport is obviously related to the crossings and anticrossings of the dressed nearly-zero-energy modes due to the coherent excitation and LZ transition between the multichains and the cavity, at which the excitation is exchanged. As a consequence the exciton transport can be controlled by the crossings and anticrossings, which depend on the configurations of the multichains. Here the polaritons act as important role. In addition, the second order partial derivative of the Hopfield coefficients of the photon is a good physical quantity to measure the extreme points of the photon occupation number and the locations of the crossings and anticrossings. The optimal parameters regions on the spaces of the TLS transition frequency detuning, the inter-chain coupling strength, and the dimerization parameter are presented. Further, the results show that the exciton transport dynamics depend on the odevity of the number and length of the chains.
The remainder of this paper is organised as follows: In Sec. II the theoretical model is introduced including the Hamiltonian of the multichain-cavity coupling system (in Sec. II.1), the formula calculating the von Neumann entropy of excitons (in Sec. II.2), and the Lindblad quantum master equation of the system with external driving field (in Sec. II.3). In Sec. II.4 the exciton transport currents and efficiency for the steady state are derived. In Sec. III the energy spectrums are given by numerically diagonalizing the Hamiltonian. In Sec. IV we present and discuss the steady-state dynamics of excitons for various configurations of the multichains, in which the effects of the cavity, the inter-chain coupling, the dimerization parameter, the TLS transition frequency detuning (in Sec. IV.1), the number and length of the chains (in Sec. IV.2) on exciton transport are analysed. Section V is a brief summary and future perspective.
Figure 1: (color online) The schematic representation of the hybrid system that the 2D TLSs array are placed in a cavity for the three configurations HO ST (a), HO TT (b), and HE ST (c). The red sphere represents the TLS with ground state $|g\rangle$ and excited state $|e\rangle$ between which the transition frequency is $\omega_{\rm a}$ and the decay rate is $\gamma$. The decay rate of the imperfect cavity is $\kappa$. The 2D TLSs array can be driven by an external coherent field with driving frequency $\omega$ via the first TLS of the first row with driving amplitude $\xi_{{\rm T},1}$ or/and via the cavity with driving amplitude $\xi_{c}$. The excitons in the multichains finally outflow from the last column into a reservoir (drain) with rate $\gamma_{\rm R}$. The intra-chain coupling strength is $h$. $\delta$ is the dimerization strength of HE chain. It exists the inter-chains couplings with the strength $\zeta$ between the nearest TLSs in two nearest chains.
II Model
II.1 Hamiltonian
The considered system is a 2D molecular aggregate placed in a single-mode cavity, which can be modelled as multichains with $L$ ($L\geqslant 2$) chains as shown in Fig. 1. The $k$th chain has $N_{k}$ identical TLSs. For simplicity, we set that the numbers of the TLSs of each chain are the same ($N_{1}=N_{2}=...=N_{L}=N$). The Hamiltonian of the multichain-cavity coupling system can be expressed as
$H_{0}=H_{\rm e}+H_{\rm c}+H_{\rm ec}$ (1)
The $H_{\rm e}$ represents the Hamiltonian of the multichains, and contains two terms as
$H_{\rm e}=\sum^{L}_{k=1}H_{{\rm e},k}+H_{\rm e,inter-chain},$ (2)
where $H_{{\rm e},k}$ are the Hamiltonian of the $k$th bare chain.
Considering the staggered hopping interaction between the nearest-neighbor TLS monomers in a chain, under the rotating frame of the driving frequency $\omega$ the Hamiltonian of the $k$th chain is described by the SSH model as
$H_{{\rm e},k}=$ $-\sum_{i=1}^{N}\Delta_{\rm a}\sigma_{k,i}^{+}\sigma_{k,i}^{-}-$
$\sum_{i=1}^{N-1}h\left[1+(-1)^{i}\delta\right](\sigma_{k,i}^{+}\sigma_{k,i+1}^{-}+{\rm H.c.}),$ (3)
where $h$ characterizes the hopping amplitude between two nearest-neighbor TLS monomers in a chain, $\delta$ represents the dimerization strength, $\sigma_{k,i}^{+}=|e_{k,i}\rangle\langle g_{k,i}|$ ($\sigma_{k,i}^{-}=|g_{k,i}\rangle\langle e_{k,i}|$) is the raising (lowering) operator of the TLS with same energy $\omega_{\rm a}$ (setting $\hbar=1$) at the $i$th site of the $k$th chain with the groud state $|g_{k,i}\rangle$ and the excited state $|e_{k,i}\rangle$ because of the identical partical, $\Delta_{\rm a}=\omega-\omega_{\rm a}$ is the TLSs transition frequency detuning relative to driving field frequency $\omega$, ${\rm H.c.}$ denotes the Hermitian conjugate. For $\delta=0$ and $\delta\neq 0$, the SSH chain behaves entirely different property, which correspond respectively homogeneous and heterogeneous coupling in a molecular chain. In addition, it exist the topologically nontrivial regime for $\delta\neq 0$. In order to emphasize the difference of the exciton transport of the two coupling cases, we designated separately them as the homogeneous (HO) and the heterogeneous (HE) coupling chain in the next content.
The inter-chain coupling term is denoted by the $H_{\rm e,inter-chain}$, whose form depends on the inter-chain coupling configuration in the 2D multichains. In this paper, we consider the two simplest types of the configuration, which are named as the square type (ST) [Fig. 1(a)] and the triangle type (TT) [Fig. 1(b,c)]. The corresponding Hamiltonian can be respectively written as
$H_{\rm e,inter-chain}=$ $\zeta\sum^{L-1}_{k=1}\sum_{i=1}^{N}(\sigma_{k,i}^{+}\sigma_{k+1,i}^{-}+{\rm H.c.})$ for ST, (4a)
$H_{\rm e,inter-chain}=$ $\zeta\sum^{P}_{k=1}\sum_{i=1}^{N-1}[\sigma_{2k-1,i}^{+}(\sigma_{2k,i}^{-}+\sigma_{2k,i+1}^{-})+\sigma_{2k-1,N}^{+}\sigma_{2k,N}^{-}+{\rm H.c.}]+$
$H_{\rm e,inter-chain}=$ $\zeta\sum^{Q}_{k=1}\sum_{i=2}^{N}[\sigma_{2k,1}^{+}\sigma_{2k+1,1}^{-}+\sigma_{2k,i}^{+}(\sigma_{2k+1,i-1}^{-}+\sigma_{2k+1,i}^{-})+{\rm H.c.}]$ for TT, (4b)
where for TT the upper bounds of $L$ in odd and even term are respectively
$P=$ $(L-1)/2$ $L$ is odd, (5a)
$P=$ $L/2$ $L$ is even, (5b)
$Q=$ $(L-1)/2$ $L$ is odd, (6a)
$Q=$ $(L-2)/2$ $L$ is even, (6b)
and $\zeta$ denotes the inter-chain coupling strength, namely the interaction strength between two nearest-neighbor TLSs in two nearest-neighbor chains. Therefore, the multichains have six possible configurations based on the permutation and combination of the inter-chain and the intra-chain coupling types, which are expressed as HO ST, HO TT, HE ST, HE TT, HOCC-HE ST, and HO-HE TT. The HO ST, HO TT, and HE ST are schematically depicted in Fig. 1(a), Fig. 1(b), and Fig. 1(c) respectively. Correspondingly, the HE TT is that the HO chains in the HO TT are substituted by the HE chains. Particularly, the HO and HE chain are arranged with staggered form in HO-HE ST and HO-HE TT, in which we set that the odd chains are HO and the even chains are HE.
In the rotating frame of the driving frequency $\omega$, the radiation part of cavity is described by $H_{\rm c}$ as
$H_{\rm c}=-\Delta_{\rm c}c^{\dagger}c,$ (7)
where $c^{\dagger}=|c\rangle\langle 0|$ ($c=|0\rangle\langle c|$) is the photon creation (annihilation) operator, the cavity mode is decided by the frequency $\omega_{\rm c}$, $\Delta_{\rm c}=\omega-\omega_{\rm c}$ is the detuning of the cavity field frequency $\omega_{\rm c}$ with respect to the $\omega$, $|c\rangle$ is the single-photon state that only the cavity field is excited, and $|0\rangle$ is the vacuum state of the exciton-cavity system. In order to satisfy the resonant condition between the cavity and the TLSs, we set $\omega_{\rm c}=\omega_{\rm a}$, hence $\Delta_{\rm a}=\Delta_{\rm c}=\Delta$.
In the rotating frame of $\omega$ the coupling between the multi-chain aggregate and the cavity is given by
$H_{\rm ec}=\sum^{L}_{k=1}\sum_{i=1}^{N}g_{k,i}(\sigma_{k,i}^{+}c+{\rm H.c.}),$ (8)
where $g_{k,i}$ describes the coupling strength between the TLS at the $i$th site in the $k$th chain and the cavity field. For the sake of simplicity, we assume that $g_{k,i}$ is uniform for all TLSs, that is $g_{k,i}=g$. Generally, $g$ depends on the collective Rabi splitting $\Omega_{\rm R}=2\sqrt{\sum_{k=1}^{L}\sum_{i=1}^{N}g^{2}_{k,i}}=2g\sqrt{LN}$.
In order to investigate the excitation dynamics, we introduce a external coherent driving in the exciton-cavity system. In this paper we choose two driving means, i.e., the driving field is exerted on the first site of the $k$th chain as well as simultaneously on this site and the cavity, which can be described respectively as
$H_{\rm d}=$ $H_{\rm dT}=\xi_{{\rm T},k}(\sigma_{k,1}^{-}e^{i\omega t}+\sigma_{k,1}^{+}e^{-i\omega t}),$ (9a)
$H_{\rm d}=$ $H_{\rm dTc}=H_{\rm dT}+\xi_{\rm c}(ce^{i\omega t}+c^{\dagger}e^{-i\omega t}),$ (9b)
where $\omega$ is the driving frequency, $\xi_{{\rm T},k}$ ($\xi_{\rm c}$) is the strength of the TLS (cavity) drive. The second mean is denoted by ”TLS+cavity” in the following. Here, we set $k=1$, namely, only the first site in the first chain is driven. The two driving forms were used to study the two-photon blockade in the cavity QED system hamsen2017two . In the rotating frame of $\omega$, the Eq. (9) can be written as
$H_{\rm d}=$ $H_{\rm dT}=\xi_{{\rm T},1}(\sigma_{1,1}^{-}+\sigma_{1,1}^{+}),$ (10a)
$H_{\rm d}=$ $H_{\rm dTc}=H_{\rm dT}+\xi_{c}(c+c^{\dagger}).$ (10b)
When the driving field is considered, the total Hamiltonian of the open quantum system is
$H=H_{0}+H_{\rm d}.$ (11)
[…]
II.2 von Neumann entropy
In order to find and distinguish accurately the locations of the energy level crossings and anticrossings of Hamiltonian (1) in the parameter space of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$, one can calculate the von Neumann entropy of the excitons by
$S(\rho_{\rm e})=-{\rm tr}(\rho_{\rm e}{\rm log}_{2}\rho_{\rm e}).$ (12)
For the hybrid system consisting of excitons and photons, its eigenstates can be indicated as $|\psi_{\rm e-p}\rangle$. Thus the reduced density matrix $\rho_{\rm e}$ of the excitons is obtained by tracing out the degrees of freedom of the photons, i.e., $\rho_{\rm e}={\rm tr}_{\rm p}(|\psi_{\rm e-p}\rangle\langle\psi_{\rm e-p}|)$. According to the conclusion of Ref. hu2018effect , the crossings (anticrossings) correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy with parameter. Therefore, the entanglement between upper and lower energy levels is maximal at the anticrossings, at which the LZ transitions occur. The limiting of distance between two anti-cross energy levels is crossing that means the energy degeneracy. At the crossings, the excitations are exchanged coherently between the multichains and the cavity. As a result, in energetically degenerate state the exciton transport can be observed tagarelli2023electrical and boosted lin2012efficient .
II.3 Transport dynamics
Under the Born-Markov approximations and the rotating-wave approximation, the dynamics of the system that exists dissipation at each TLS in the multichains, the coupling between the first site of the first chain and the driving field as well as between the last site of each chain and the reservoir, are governed by the Lindblad quantum master equation lindblad1976generators . Then the exciton transport dynamics in the system can be investigated by the master equation
$\dot{\rho}=i[\rho,H]+\sum^{L}_{k=1}(D_{{\rm d},k}[\rho(t)]+D_{{\rm R},k}[\rho(t)]),$ (13)
where the first term in the righthand side accounts for the free evolution under the total Hamiltonian (11), the second term characterizes the decay of excitons due to the spontaneous emission with
$D_{{\rm d},k}[\rho]=\sum^{N}_{i=1}\gamma\mathcal{L}_{\sigma^{-}_{k,i}}[\rho]+\kappa\mathcal{L}_{\rm c}[\rho],$ (14)
where $\gamma$ and $\kappa$ are the decay rates of the TLSs and the cavity photons respectively, and $\mathcal{L}_{x}[\rho]=x\rho x^{\dagger}-\dfrac{1}{2}\left\{\rho,x^{\dagger}x\right\}$ is the Lindblad superoperators with standard form. The term $D_{{\rm R},k}[\rho(t)]$ in Eq. (13) represents the exciton exchange between the last site of the $k$th chain and the exciton drain, which is given by
$D_{{\rm R},k}[\rho]=\gamma_{\rm R}\bar{n}_{\rm R}\mathcal{L}_{\sigma_{k,N}^{+}}[\rho]+\gamma_{\rm R}(\bar{n}_{\rm R}+1)\mathcal{L}_{\sigma_{k,N}^{-}}[\rho],$ (15)
where $\gamma_{\rm R}$ is the exciton exchanging rate between the last site of each chain and the drain, $\bar{n}_{\rm R}$ is the average exciton number in the drain. The last sites of each chain exist coupling with the drain, so $k\in\{1,2,...,L\}$. In Eq. (15), the two terms denote respectively the inflow and outflow of excitons from the last column TLSs of the multichains to the drain. In order to show the net transportation of exciton, we set as $\bar{n}_{\rm R}=0$ which is reasonable as the photosynthesis process in nature. Therefore, the exciton is unidirectionally transferred from the last column of the multichains to the drain without exciton reflection.
II.4 Efficiency of exciton transport
Figure 2: (color online) The energy spectra of the exciton-cavity coupling system as a function of the inter-chain coupling strength $\zeta$ for different configurations HO ST (a), HE ST (b), HO-HE ST (c), with the colormap representing the Hopfield coefficients of the content of cavity photons by solving the eigenvalues and eigenstates of the Hamiltonian (1) in zero- and single-excitation subspace. (d-f) The second order partial derivative of the Hopfield coefficients of the content of cavity photons with respect to $\zeta$, which contain the discontinuity and the extremum corresponding to the marked locations existing evident changes in steady-state dynamics. (g-i) The corresponding von Neumann entropy $S$ of excitons by calculating Eq. (12). The color of the lines has not means just to distinguish each other in (d-i). The cavity-dressed energy level crossings (anticrossings) producing obvious effect on exciton transport are marked by the black (violet) vertical lines and the data texts of their locations, which correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy. The corresponding parameters are $L=2$, $N=8$, $\Delta_{\rm a}=\Delta_{\rm c}=0$ eV, $h=1$ eV, $\delta=0.5$, and $\Omega_{R}=1$ eV.
Figure 3: (color online) Similar as Fig 2, but for different configurations HO TT (a, d, and g), HE TT (b, e, and h), and HO-HE TT (c, f, and i).
Figure 4: (color online) Similar as Fig 2, but for even chain $N=8$ and for different configurations HE ST (a, e, and i), HE TT (b, f, and j), HO-HE ST (c, g, and k), and HO-HE TT (d, h, and l) with dimerization parameter $\delta$. The parameter used is $\zeta=0.1$ eV, the others are the same as those in Fig. 2.
Figure 5: (color online) Similar as Fig 4, but for odd chain $N=9$.
To show the exciton transport dynamics of the system containing a 2D multichains and a cavity, the exciton transport efficiency is defined and derived in this subsection. Generally, the exciton transport efficiency is characterized by the ratio of the outgoing $(I^{\rm s}_{\rm o})$ and input $(I^{\rm s}_{\rm i})$ exciton currents for steady state liu2019vibration ; wei2019enhanced , that is
$\eta=\lim_{t\to\infty}\frac{I_{\rm o}}{I_{\rm i}}=\frac{I_{\rm o}^{\rm s}}{I_{\rm i}^{\rm s}},$ (16)
where the superscript ”s” refers to the steady state. On the other hand, the exciton occupation number can display the transport machanism and be used to solve the $I_{\rm i}$ and $I_{\rm o}$. The exciton occupation number on the $j$th site of the $k$th chain is given by
$p_{k,j}(t)\equiv{\rm Tr}\{\rho\sigma^{+}_{k,j}\sigma^{-}_{k,j}\}.$ (17)
Then combining Eq. (13) with Eq. (17), the equation that $p_{k,j}(t)$ obeys is derived as
$\dot{p}_{k,j}(t)=$ $i{\rm Tr}\{[\rho,H]\sigma^{+}_{k,j}\sigma^{-}_{k,j}\}+$
${\rm Tr}\{\sum_{\mu={\rm d,R}}\sum^{L}_{k=1}D_{\mu,k}[\rho]\sigma^{+}_{k,j}\sigma^{-}_{k,j}\},$ (18)
where the trace is operated throughout the Hilbert space of the total system. The lefthand side of Eq. (18) means the rate of the change of $p_{k,j}(t)$ with time, thus the first term (the second term) of the righthand side represents the input (outgoing) exciton current at the $j$th site in the $k$th chain. According to the explanation, under the TLS and TLS+cavity drive the input exciton current of the multichains can be written respectively as
$I_{\rm i}(t)=$ $i{\rm Tr}\{[\rho,H_{\rm dT}]\sigma^{+}_{1,1}\sigma^{-}_{1,1}\},$ (19a)
$I_{\rm i}(t)=$ $i{\rm Tr}\{[\rho,H_{\rm dT}]\sigma^{+}_{1,1}\sigma^{-}_{1,1}+[\rho,\xi_{c}(c+c^{\dagger})]c^{\dagger}c\},$ (19b)
which is related to the exciton injection at the first site in the $k$th chain or/and in cavity via the external coherent field per unit time. Substituting Eq. (9) for $H_{\rm dc}$ and $H_{\rm dTc}$ in Eq. (19), under the zero- and single- excitation subspace, the input exciton currents are simplified to
$I_{\rm i}(t)=$ $i\xi_{{\rm T},1}(\langle 1,1|\rho|0\rangle-{\rm H.c.}),$ (20a)
$I_{\rm i}(t)=$ $i\xi_{{\rm T},1}(\langle 1,1|\rho|0\rangle-{\rm H.c.})+i\xi_{\rm c}(\langle c|\rho|0\rangle-{\rm H.c.}).$ (20b)
According to the Hermitian property of density matrix, under the TLS and TLS+cavity drive the input exciton currents are finally written as
$I_{\rm i}(t)=$ $2\xi_{{\rm T},1}{\rm Im}[\langle 0|\rho|1,1\rangle],$ (21a)
$I_{\rm i}(t)=$ $2\xi_{{\rm T},1}{\rm Im}[\langle 0|\rho|1,1\rangle]+2\xi_{c}{\rm Im}[\langle 0|\rho|c\rangle].$ (21b)
Analogous to the input exciton current, the outgoing exciton current can be defined as the outflow of exciton transfering from the last column in the multichains to the drain per unit time. According to the means of the third term of righthand in Eq. (18), it can be given by
$I_{\rm o}(t)={\rm Tr}[\sum^{L}_{k=1}D_{{\rm R},k}[\rho(t)]\sigma_{k,N}^{+}\sigma_{k,N}^{-}],$ (22)
where each chain has exciton outgoing to the drain for all configurations, thus $k\in\{1,2,...,L\}$. Pluging the Eq. (15) into the Eq. (22), under the zero- and single- excitation assumption, the form of the outgoing exciton current is simplified as
$I_{\rm o}(t)=\gamma_{\rm R}\sum^{L}_{k=1}p_{k,N},$ (23)
where $p_{k,N}$ is the occupation number on the last site of the $k$th chain.
III Spectrum of coupling system
There are a large number of studies on the spectra of the SSH Hamiltonian for the open boundary conditions sirker2014boundary ; nie2020bandgap ; wei2022cavity ; cai2019observation . The topological phases of the SSH chain depend on the parity of the number of an 1D TLSs chain and the qubit-qubit coupling configuration. According to the Ref. cai2019observation , the dimerization parameter $\delta>0$ ($\delta<0$) corresponds to the topological winding number $\nu=1$ ($\nu=0$), i.e., the topologically nontrivial (trivial) regime. In Ref. sirker2014boundary , the spectra show that there is one zero-energy mode in whole range of $\delta$ for odd and are two localized nearly-zero-energy modes in the topologically nontrivial regime $\delta>0$ for even chain. The finite gap between the two edge modes leads to the formation of the hybrid edge modes that overlap with each other, generating the polaritonic states when the edge modes are dressed by cavity due to the coupling between the excitons and the photons. Comparing to the single chain system, the presence of the inter-chain coupling induces the extension of the spectra of the 2D multi-chain system. As we all known, the gap corresponds to the energy level anticrossings, at which the LZ transitions among the levels occur in multi-level LZ system, adding greatly to the complexity of the system dynamics hu2018effect . What’s more, the energy level crossings mean that the corresponding quantum states are degeneracy. In the exciton-cavity coupling system, at the crossings the excitations are coherently exchanged between the TLSs and the cavity. Therefore, the crossings and anticrossings of the edge modes dressed by the cavity can lead to the abundant exchange of excitation between the multichains and the cavity, increasing or decreasing polaritons. The forming polaritons act as intermediary in exciton transport, where the excitons are transported among site $(k,1)~{}\rightarrow$ cavity $\rightarrow$ site $(k,N)$ without experiencing the dissipation of the other TLSs on multichains wei2022cavity . Therefore, the polaritons lead to the exciton transport efficiency and currents increasing, which can be controlled by the crossings and anticrossings.
Figure 6: (color online) The dependence of the steady-state occupation numbers (a1 and b1) $p_{11}$ on site 11, (a2 and b2) $p_{\rm bulk}$ on bulk sites, (a3 and b3) input exciton current $I_{\rm i}$, (a4 and b4) outgoing exciton current $I_{\rm o}$, and (a5 and b5) transport efficiency $\eta$ on the inter-chain coupling strength $\zeta$ (a1-5) and the dimerization parameter $\delta$ (b1-5) for different configurations without cavity. In (b1-5) the solid and dash lines correspond respectively to $N=8$ and 9. The parameters used are $\zeta=0.1$ eV, $\xi_{{\rm T},1}=\xi_{\rm c}=0.1$ eV, $\gamma_{\rm d}=0.01$ eV, $\gamma_{\rm R}=0.1$ eV, and the others are the same as those in Fig. 2.
Figure 7: (color online) The dependence of the steady-state dynamics of the hybrid system on the inter-chain coupling strength $\zeta$ in the presence of cavity under TLS drive (a1-6) and TLS+cavity drive (b1-6). (a5 and b5) The occupation number of photon with $\zeta$, in which the extreme points are marked by the vertical fine lines, and the others are similar to Fig. 6. The parameter used is $\kappa=0.01$ eV, and the others are the same as those in Fig. 6.
Figure 8: (color online) The dependence of the steady-state dynamics of the hybrid system on the dimerization parameter $\delta$, and the others are similar to Fig. 7. The parameters used are the same as those in Fig. 7.
In order to analyse the spectra of the coupling system consisting of the multichains and the cavity, its eigenvalues can be obtained by numerically diagonalizing the Hamiltonian (1). Further, the formation of polaritons can be investigated by calculating the Hopfield coefficients $|\alpha_{c}|^{2}$. When the $j$th normalized eigenvector is expanded as $|\psi_{j}\rangle=\Sigma_{k}^{L}\Sigma_{i}^{N}\alpha_{k,i}|k,i\rangle+\alpha_{c}|c\rangle$, $|\alpha_{k,i;c}|^{2}$ denotes Hopfield coefficient, which characterizes the corresponding hybridizing weight of different excitonic and photonic contents. The color maps Fig. 2 (a-c), 3 (a-c) and Fig. 4 (a-d), 5 (a-d) show respectively the spectra of the exciton-cavity coupling system as functions of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$ for different coupling configurations, and contain the photonic Hopfield coefficient $|\alpha_{c}|^{2}$. One can see that the zero-energy modes exist in the spectra for each situation. $|\alpha_{c}|^{2}\neq 0$ means that the energy level is dressed by the cavity mode. It is obvious that only the nearly-zero-energy modes are distinctly dressed by the cavity mode, generating corresponding polaritonic branches, because the high excited states are hardly dressed for the given coupling strength between excitons and photons. In addition, there are the crossings (anticrossings) in the spectra, in which the crossings (anticrossings) between the nearly-zero-energy modes dressed by the cavity mode are our focus. It can be seen that the ratio of the distributed photonic contents on the upper and lower polaritonic branches vary with $\zeta$ and $\delta$. It means that the coherent exchange of excitation or the LZ transition of photons occurs, inducing the change of the exciton transport efficiency and currents, which will be analysed in the next section by the steady-state dynamics. Further, it can be discovered that the number and locations of crossings (anticrossings) affecting transport depend on the coupling configurations of TLSs in the range of $\zeta$ and $\delta$. In comparison, the effect of $\zeta$ on the crossings (anticrossings) is more significant than that of $\delta$.
In order to find the changing regularity of the photonic Hopfield coefficient $|\alpha_{c}|^{2}$, we plot the scatter maps of the second order partial derivatives $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of $|\alpha_{c}|^{2}$ with respect to $\zeta$ in Fig. 2 (d-f), 3 (d-f) and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with respect to $\delta$ in Fig. 4 (e-i), 5 (e-i) that are related to the obvious changes of the steady-state dynamics of exciton transport. To determine and distinguish accurately the locations of the crossings and anticrossings in spectra, the corresponding von Neumann entropy $S$ of excitons is calculated through Eq. (12), and is shown in Fig. 2 (g-i), 3 (g-i), 4 (i-l), and 5 (i-l). The crossings (anticrossings) determined by $S$ is signed in the figure of the spectra, which correspond to the saltatorial intersections (maximum values or acclive intersections) of $S$. Significantly, we discovery that the incontinuity locations of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with $\zeta$ and $\delta$ correspond to the crossings, and the extreme points of that correspond to the anticrossings. As shown in Fig. 2(d), the green and red lines of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ are discontinuity at $\zeta=-0.3$ where the dressed energy levels exist the crossing [Fig. 2(a) and 2(g)], and the green and cyan lines of that appear the maximal values as well as the blue and green lines of that appear the minimal values at $\zeta=-0.21$, -0.02 respectively where the dressed energy levels exist the anticrossings [Fig. 2(a) and 2(g)]. We also can see that the green, blue, and cyan lines are discontinuity at $\zeta=0.07$, 0.09, 0.22 respectively [Fig. 2(e)] corresponding to the crossings of the dressed energy levels [Fig. 2(b) and 2(h)]. The green and magenta lines behave as maximal values, and the magenta and cyan lines behave as minimal values at $\zeta=-0.03$, 0.48 respectively [Fig. 2(e)] corresponding to the anticrossings of the dressed energy levels [Fig. 2(b) and 2(h)]. Similarly, the maximal values occur at $\zeta=-0.1$, -0.05, and 0.03 in the black, green, and blue lines respectively, and the minimal values occur at $\zeta=-0.05$, 0.03 in the red and cyan lines respectively [Fig. 2(f)], which correspond to the anticrossings of the dressed energy levels [Fig. 2(c) and 2(i)]. Therefore, the second order partial derivative of the photonic Hopfield coefficients is a good physical quantity to determine the crossings and anticrossings of spectra. Comparing the spectra under different configurations, one can discover that the gap in the range $\zeta\in[-0.21,-0.02]$ for the configuration HO ST and its range [Fig. 2(a)] is respectively widened and extended by the heterogeneous coupling in multichain as shown in Fig. 2(b). For the mixed configuration HO-HE ST new gaps appear near $\zeta=0$ as shown in Fig. 2(c). About $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ in Fig. 3 (d-f), 4 (e-i), and 5 (e-i), the conclusion is similar as that of Fig. 2 (d-f), thus it is no longer stated hereinbelow. When the inter-chain coupling is TT, the spectra is shown in Fig. 3(a-c). It can be seen that the gaps near $\zeta=0$ narrow, and its ranges in $\zeta$ reduce compared with Fig. 2(a-c). In the space of $\delta$, the spectra are presented in Fig. 4 (a-d) for even chain and Fig. 5 (a-d) for odd chain. In Fig. 4 (a-d) we can see that the configuration HE ST has more crossings of and energy levels of the dressed nearly-zero-energy modes than HE TT. In mixed multichains the presence of the HO chain leads also to the decrease of the crossings and energy levels, comparing Fig. 4 (a) and 4 (c). However, the crossings at $\delta=0.3$ and 0.36 [Fig. 4 (c)] covert to anticrossing at $\delta=0.19$ [Fig. 4 (d)] because of TT coupling. The energy levels of the dressed nearly-zero-energy modes cross for even chain and avoid crossing for odd chain. Therefore, there is less crossings affecting exciton transport for odd chain [Fig. 5 (a-d)].
In the next section, the steady-state dynamics with $\zeta$ and $\delta$ will be shown, in which the relation between the dressed crossings (anticrossings) and the occupation number of photons can be discussed. The changes of exciton transport currents and efficiency can be explained. We aim at revealing the mechanism of exciton transport by finding the optimal range of parameters on exciton transport and comparing the spectra and steady-state dynamics for different configurations.
Figure 9: (color online) The dependence of the steady-state dynamics on the inter-chain coupling strength $\zeta$ and the detuning $\Delta$ of the driving frequency relative to the TLS without cavity under TLS drive (a-f, ab-fb), and with cavity under TLS drive (g-l, gc-lc) and under TLS+cavity (the TLS and cavity drive exist simultaneously) drive (m-r) for different configurations. The corresponding physical quantities of each row are shown at right. The zones that cavity produces significant effect on exciton transport efficiency are marked by ellipses in (g-l), the corresponding zones are marked in (ab-fb) and (gc-lc). The parameters used are the same as those in Fig. 7.
Figure 10: (color online) The dependence of the steady-state transport efficiency $\eta$ on the dimerization parameter $\delta$ and the detuning $\Delta$ of the driving frequency relative to the TLS, in which the corresponding conditions are written at the top and the right of this figure. The parameters used are the same as those in Fig. 7.
Figure 11: (color online) The steady-state dynamics as a function of the number $L$ of the chains without a1-g1) and with (a2-h2) cavity for different configurations, in which $p_{1,i}$, $p_{L,i}$, and $p_{\rm bulk-in}$ represent respectively the occupation number of the first row, the last row, and the bulk state denoted as ”bulk-in” containing all TLSs inside of the multichains. The parameter used is $N=10$, and the others are the same as those in Fig. 7.
Figure 12: (color online) The steady-state dynamics as a function of the length $N$ of chains for different configurations and number of the chains, (a-d) without cavity under TLS drive, (e-h) with cavity under TLS drive. The parameters used are the same as those in Fig. 7.
IV The exciton transport dynamics by multichains
In this section, we discuss the effects of the inter-chain coupling, the dimerization parameter, the number and length of the chain on the exciton transport dynamics by numerically solving the master equation Eq. (13) under an external driving, which are displayed in two subsections below. According to the analysis in the Sec. III, the nearly-zero-energy modes are dressed by the cavity for strong exciton-cavity coupling, namely the polaritonic branches forming. In order to let the polaritons act as a dominated role in exciton transport, we set that the Rabi splitting is far stronger than the decay rates of the cavity and the TLSs, i.e., $\Omega_{R,k}\gg(\kappa,\gamma)$.
IV.1 The effects of the inter-chain coupling and the dimerization parameter
Here, we investigate primarily the corresponding relation between the energy level crossings (anticrossings) dressed by the cavity and the exciton transport dynamics to explain the exciton transport mechanism. Based on the relation and the properties of exciton transport dynamics, the optimal parameter ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$, as well as the optimal coupling configurations are expected to be presented.
Fig. 6 shows the dependence of the steady-state exciton transport dynamics of the multi-chain system without cavity on $\zeta$ and $\delta$ for the six chosen configurations. From the Eq. (23), it can be known that the outgoing exciton current $I_{\rm o}$ is directly proportional to the total occupation number $\Sigma_{1}^{L}p_{k,N}$ of the outgoing sites, thus only $I_{\rm o}$ is shown in following figures, and $\Sigma_{1}^{L}p_{k,N}$ is denoted by $p_{N}$. One can see that the effects of the inter-chain coupling and the dimerization parameter on exciton transport are significantly different. The steady-state occupation number $p_{11}$ of site ”11” (The two ”1” denote respectively the 1st row and the 1st column in the array of TLSs shown in Fig. 1. This representation is similar below, unless otherwise noted.), $p_{\rm bulk}$ of bulk sites (The ”bulk” contains the all sites from the 2nd to the $(N-1)$th column), the input and outgoing exciton current $I_{\rm i,o}$, and the exciton transport efficiency $\eta$ with $\zeta$ are symmetric about $\zeta=0$ in $[-1,1]$, but with $\delta$ are not.
For the six configurations, the descending sort of efficiency is successively HO-HE ST, HO ST, HO TT, HO-HE TT, HE ST, HE TT in $\zeta\in[-1,1]$ [Fig. 6(a5)]. For the HO-HE ST, near $\zeta=0.3$ the sites of bulk [Fig. 6(a2)], input, and outgoing possess commonly relatively high occupation numbers of exciton, which support exciton transport in multichains. Therefore, the exciton transport efficiency, input and outgoing currents are the largest near $\zeta=0.3$ for the HO-HE ST [Fig. 6(a3,4,5)]. On the contrary, for the HO-HE ST at $\zeta=0$ and $\zeta\rightarrow\pm 1$, although the site 11 possesses more excitations [Fig. 6(a1)], the exciton occupation number $p_{N}$ at outgoing sites is too few to transfer to the drain, leading the low efficiency and exciton currents. Specially, the excitons are distributed mainly at bulk sites for $\zeta\rightarrow\pm 1$ [see cyan line in Fig. 6(a2)], which is beneficial to energy storage and can be used to design quantum battery quach2022superabsorption . Except for the HE ST, the other configurations also have the property for $\zeta\rightarrow\pm 1$, as well as HO TT and HO-HE TT also do near $\zeta=\pm 0.4$. However, for HO ST near $\zeta=\pm 0.35$ and ±$\pm 0.97$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ are relatively high, thus the ability both of energy transport and storage are strong. Near $\zeta=\pm 0.7$ for HO ST, although the exciton transport efficiency is large, exciton currents are relatively small. Thus it is not beneficial to the exciton transport.
It has four configurations HE ST, HE TT, HO-HE ST, HO-HE TT containing HE chain. Because of the different spectra between odd and even HE chain, in Fig. 6(b1-5) there are two situations $N=8,~{}9$ about the steady-state dynamics with $\delta$ in the absence of cavity. It is obvious that the exciton currents and transport efficiency are the largest near $\delta=0.3$ for the configurations HE ST and HO-HE ST of $N=8$ than that in the other range of $\delta$ and for the other configurations [Fig. 6(b3-5)], which are good for effective exciton transport. However, the HE TT of $N=9$ is the most beneficial to energy storage because of the high occupation number $p_{\rm bulk}$ of bulk sites [Fig. 6(b2)] as well as the lowest exciton currents and transport efficiency. Particularly, the exciton transport efficiency is high for $\delta<0$, but the exciton currents approach to zero, as shown in Fig. 6(b3-5), thus there is no real exciton transport. The reason is that the site 11 can not be coherently excited by drive, namely the excitons can not be injected in multichains [Fig. 6(b1,2)]. When $\delta>0.35$, the exciton currents and transport efficiency decrease sharply with $\delta$ or too small, because the excitons are localized at the left-edge site so that the occupation number of the right-edge sites is small for the HO-HE ST and HO-HE TT of $N=8,9$ or approaches to zero for the others.
In order to investigate the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ on the exciton transport dynamics in the presence of cavity, we plot the Fig. 7 and 8 corresponding to Fig. 6, in which the occupation number $p_{\rm c}$ of photons is added. We pay closely attention to the change of the steady-state dynamics with $\zeta$ and $\delta$ at the crossings (anticrossings) of nearly-zero-energy modes dressed by cavity.
Under the TLS and the TLS+cavity drive, the effect of the inter-chain coupling $\zeta$ on the steady-state dynamics of the multi-chain system in a cavity is presented for different configurations in Fig. 7(a1-6) and 7(b1-6) respectively. Particularly, to analyse what role the cavity acts as in exciton transport, the dependence of the occupation number $p_{\rm c}$ of photons on $\zeta$ is shown, and its extreme points are marked by the vertical fine lines in Fig. 7(a5), which correspond to the locations of the crossings and anticrossings in spectra, as well as of the incontinuity and extreme points of the second order partial derivative $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of the photonic Hopfield coefficients, marked by the vertical fine lines in Fig. 2 and 3. In consequence, the second order partial derivative of the photonic Hopfield coefficients can also describe the change of the occupation number of photons. It can be seen that the presence of the cavity breaks the symmetry of $I_{\rm i,o}$ with respect to $\zeta=0$ for configurations HO ST, HE ST, HE TT, and HO-HE ST [Fig. 7(a3,4)]. Comparing with Fig. 6, one can discover that the exciton transport is suppressed by the cavity in the range of $\zeta\neq 0$. For example, near $\zeta=-0.3$ $p_{\rm c}$ is high [Fig. 7(a5)], suppressing the exciton currents in the main channel with the multichains for HO ST [Fig. 7(a3,4)]. On the other hand, there are three configurations HE ST, HE TT, and HO-HE TT for which the cavity induces significant increase of $I_{\rm i,o}$ and $\eta$ near $\zeta=0$. In the situation, the exciton transport is assisted by polaritons, in which $p_{\rm c}$ is high, leading the increase of exciton occupation number $p_{N}$ on the last column sites [Fig. 7(a4)]. In fact, $\zeta=0$ means that only one chain undertakes the exciton transport, in which polariton-assisted exciton transport is demonstrated wei2022cavity . Because of few excitations injected [Fig. 7(a1)] for HO ST, HO TT and few $p_{N}$ [Fig. 7(a4)] for HO-HE ST near $\zeta=0$, the cavity has no effect on exciton transport. What’s more, it can be seen obviously that the peaks of $p_{\rm c}$ and of $I_{\rm i,o}$ are at different locations. The reason is that at the peak of $p_{\rm c}$ the multichains do not occupy enough excitations when it dominates the exciton transport, causing relatively low $I_{\rm i,o}$. As shown in Fig. 7(b5), adding the cavity drive $p_{\rm c}$ increase largely, but $I_{\rm i,o}$ and $\eta$ do not. Near $\zeta=0$, because of the transfer of excitations the occupation number $p_{11}$ of site 11 increases [Fig. 7(b1)], causing the decrease of $p_{\rm c}$ for HE ST, HE TT, HO-HE ST, and HO-HE TT [Fig. 7(b5)], suppressing the exciton transport of the main channel with the porlaritons [Fig. 7(b3,4,6)]. However, one can see by comparing Fig. 7(b3,4) and Fig. 7(a3,4) that the increase of $p_{\rm c}$ suppresses significantly the exciton currents of the main channel with the multichains near $\zeta=0.2$ for HO-HE ST and near $\zeta=0.35$ for HO ST. Therefore, the exciton currents $I_{\rm i,o}$ and transport efficiency $\eta$ can be adjusted by the cavity via photons. According to different configurations, we have the conclusion that the coupling type ST is more favorable to exciton transport than TT, and the optimal locations of exciton transport for HO and HE are different in parameter space of $\zeta$, which are consistent with the spectra. In a word, the competition between cavity photons and excitons is the key affecting exciton transport.
Fig. 8(a1-6) and 8(b1-6) show the dependence of the steady-state dynamics on the dimerization parameter $\delta$ under the TLS and the TLS+cavity drive in the presence of cavity respectively. We analyse firstly the effects of $\delta$ and cavity on exciton transport under the TLS drive. Although the exciton transport efficiencies are centered approximatively on $\delta=0$ [Fig. 8(a6)], there are no the exciton currents in $\delta<0$ except for the configurations HE ST and HO-HE ST of $N=9$ [Fig. 8(a3,4)]. The root cause is that the excitation can not enter to multichains through the site 11 [Fig. 8(a1,3)]. It can be seen that for the configurations HE ST and HO-HE ST of $N=8$ in the range $0.350<\delta<$ 0.35 the presence of cavity suppresses significantly the exciton currents $I_{\rm i,o}$ [Fig. 8(a3,4)] and transport efficiency $\eta$ [Fig. 8 (a6)] by comparing with Fig. 6(b3-5). Particularly, at $\delta=0.26$ for HE ST and $\delta=0.36$ for HO-HE ST of $N=8$, the troughs of $I_{\rm i,o}$ and $\eta$ appear, which correspond to the peak and the trough of $p_{\rm c}$ [Fig. 8(a5)] and to the crossings of the dressed near zero-energy modes marked by the vertical fine lines in Fig. 4, respectively. The reason for the former is that the excitation on the first and the last column TLSs is transferred to cavity through coherent exciting at the dressed crossings as shown in Fig. 8(a1,3,4), suppressing the exciton currents and transport efficiency on the main channel with multichains. But for the latter, the occupation number $p_{N}$ of exciton on the last column TLSs is too few, limiting the exciton transport. At $\delta>0.26$ for HE ST and $\delta>0.36$ for HO-HE ST of $N=8$, the limited photon number $p_{\rm c}$ [Fig. 8(a5)] and the decreased occupation number $p_{\rm bulk}$ on bulk sites [Fig. 8 (a2)] lead to that the polariton becomes the main channel, thus the exciton transport is strengthened by the cavity comparing with Fig. 6(b3,4). However, for HO-HE TT both the multichains and the polariton dominate the exciton transport because of the increase of $p_{\rm bulk}$ and $p_{\rm c}$ with $\delta$. For the odd chains, in about $\delta>0.35$ the low $p_{N}$ is main reason limiting exciton transport. When the cavity drive is added based on the TLS drive, there are distinct effects on exciton transport in $0.350<\delta<$ 0.35 for HE ST and HO-HE ST of $N=8$. The increased $p_{\rm c}$ [Fig. 8(b5)] induces the increase of polaritons, enhancing the exciton transport [Fig. 8(b3,4)] than that under TLS drive [Fig. 8(a3,4)], but in $\delta>0.35$ because of the low $p_{\rm c}$ and $p_{\rm bulk}$ [Fig. 8(b2)] the exciton transport is limited for HE ST of even chain. On the contrary, because of the high $p_{\rm c}$ and $p_{\rm bulk}$ the exciton transport is heightened for HO-HE ST and HO-HE TT of odd chain. Nevertheless, the high $p_{\rm bulk}$ and $p_{\rm c}$ make that the exciton currents and transport efficiency still are high in $\delta>0.7$ for HO-HE ST compared with that under TLS drive. Interestingly, the exciton currents and transport efficiency are significantly large due to the cavity drive in $\delta<0$ for all configurations of $N=9$, which break the universal situations that the exciton currents and transport efficiency are higher in $\delta>0$ than that in $\delta<0$. Specially, when the cavity is driven, the occupation number $p_{11}$ of site 11 is no longer necessary to transport efficiently exciton [Fig. 8(b1)]. In the condition, the high $p_{\rm bulk}$ [Fig. 8(b2)] and $p_{\rm c}$ [Fig. 8(b5)] illustrate that both the multichains and the polaritons participate the exciton transport. Although $p_{11}$ no longer is the condition restricting exciton transport under the TLS+cavity drive, $p_{N}$ still is. Therefore, the low $p_{N}$ limits the exciton transport for HE TT and HO-HE TT of $N=9$ in $\delta<0$. To sum up, in $\delta>0$ the cavity have significant influence to exciton transport, which the exciton transport is suppressed in weak dimerization strength and is enhanced in strong dimerization strength for even chain under TLS drive. In the space of $\delta$, the coupling type ST on even chain is more beneficial to exciton transport. However, when the cavity drive is exerted, ST on odd chain is good for exciton transport.
When the TLS transition frequency detuning $\Delta$ is considered, we plot the steady-state dynamics of the hybrid system in Fig. 9 and 10, which show respectively the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ for different configurations. Comparing with Fig. 2 - 5, the structures of exciton transport efficiency are similar to those of spectra. In Fig. 9 the zones of $\zeta$ and $\Delta$ that cavity has obviously effect on exciton transport are marked in Fig. 9(ab-fb, g-l, and gc-lc) by ellipses through comparing to Fig. 9(a-f). According to the effect these configurations can be classified into two categories. One is HO ST and HO TT that the exciton transport is suppressed, the other is the remaining configurations that the exciton transport is enhanced. For the former the transfer of excitation to the cavity leads to the exciton transport channel with the multichains suppressed [Fig. 9(gc,hc)]. For the letter, as shown in Fig. 9(cb-fb) $p_{\rm bulk}$ is too small in the zone existing exciton transport in the absence of the cavity, thus the polaritons dominate the transport of exciton, in which the occupation number $p_{\rm c}$ of the cavity photons is relatively large [Fig. 9(ic-lc)]. Further, we can illustrate the mechanism through the spectra. Interestingly, the structures of $p_{\rm c}$ [Fig. 9(gc-lc)] are similar to those of the cavity-dressed energy levels [Fig. 2 and 3]. In the intensified zone by the cavity there are the cavity-dressed anticrossings [Fig. 2 and 3], thus LZ transition is the foundation that the cavity effects exciton transport. Particularly, in the high efficient zone of exciton transport there are two anti-crossed zonate zones with large $p_{\rm c}$ [Fig. 9(ic)] corresponding to the dressed anticrossings for HE ST. When the cavity drive is exerted further, for the two categories of configurations the exciton transport is further suppressed and enhanced by the cavity respectively. For HE ST the effect of cavity drive on exciton transport is the biggest than the other configurations, in which the high efficient zone is significantly extended. In parameter space of $\delta$ and $\Delta$, as shown in Fig. 10 the steady-state dynamics depend on the coupling configuration, the cavity, and the method of drive. For HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ the cavity has apparent effects on exciton transport efficiency in the zone of the fully dimerized limit, namely, $\delta\rightarrow\pm 1$, and near the zero detuning [Fig. 10(a-d, g-i, m-p)]. Nevertheless, the effect of the cavity is suppressed for HO-HE TT [Fig. 10(s-v)] because the HO chain and the diagonal inter-chain coupling suppress the dimerized effect. Undoubtedly, in the zone enhancing exciton transport the polaritons act as prime role whose mechanism is coherent excitation (LZ transition) located at the cavity-dressed crossings (anticrossings) according to the analysis on Fig. 4, 5 and 8. When the cavity drive is added, it can be seen that the high efficient zones are extended and the exciton transport efficiency is strengthened in the zone of the fully dimerized limit and near the zero detuning for HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ [Fig. 10(e, f, k, l, q, r)]. If the concern is only on the optimal efficiency of exciton transport, HE ST is the preferred configuration in the parameter space of $\zeta$ and $\Delta$ under the TLS drive in the presence of the cavity, as well as HE TT is the preferred configuration in the parameter space of $\delta$ and $\Delta$ under the TLS+cavity drive. Significantly, according to the results the exciton transport can be controlled by the coupling configurations, the cavity, the drive methods, the inter-chain coupling, the dimerization parameter, and the TLS transition frequency detuning.
IV.2 The effects of the number and the length of chains
In this part, we discuss the dependence of exciton transport dynamics on the number $L$ and length $N$ of the chains, and contribute to reveal the inter-chain joint effect, to find optimal $L$, $N$, and configurations with high transport efficiency and currents in Fig. 11 and 12. It can be seen in Fig. 11(c1111, g1111, d1111) that in the absence of the cavity the number $L$ of chains has obvious effect on exciton transport of the three configurations SSH ST, TC-SSH ST, and TC-SSH TT for $L\leq 6$. For HE TT, although the input exciton current is large, the occupation number $p_{k,N}$ of the last column limits the outgoing exciton current, leading to that $L$ has not effect on exciton transport. In fact, because of the inefficient transport of excitons among chains for HE TT, the most of excitons locate at the upper edge sites [Fig. 11(b1)], which is similar to skin effect. On the contrary, because of the diffusion of excitons among multichains the inner occupation number $p_{\rm bulk-in}$ increase and the occupation number of the edge chains decrease with $L$ for HE ST as shown in Fig. 11(b1, e1, f1). Therefore, for HE ST the excitons are aggregated on the inside of the multichains with large $L$ [Fig. 11(e1)], thus it is perfect candidate applied to quantum battery in the six configurations. In the presence of the cavity the steady-state dynamics represent abundant phenomena as shown in Fig. 11(a2-h2). The prominent one is that the exciton currents and transport efficiency present distinct periodic oscillation with $L$ for HE TT and HO-HE TT, in which the peaks (troughs) of $I_{\rm i,o}$ and $\eta$ correspond to the odd (even) numbers of $L$ for $N=10$ [Fig. 11(c2, g2, d2)]. The result is contrary to that of single chain. The difference between the steady-state dynamics of the odd and of the even number becomes small with $L$. In addition to, the exciton transport is strengthened by the cavity for HE ST and HE TT.
There is an obvious character that the exciton transport efficiency $\eta$ behaves oscillation with the length $N$ of the chains in Fig. 12. We firstly discuss the situation without cavity in Fig. 12(a-d). The oscillation is weak for HO ST relative to the other configurations and depends on the parity of $L$ and $N$ for HO TT in the absence of cavity. Thus the diagonal inter-chain coupling induces difference between exciton transports of odd and of even $N$ in long HO TT multichains as shown in Fig. 12(b, d). It can be found that because of the dimerization the prominent oscillation with $N$ exists in the short multichains, but disappears in the long multichains for HE ST and HE TT [Fig. 12(a-d)]. In fact, the dimerization hinders the exciton diffusing on the multichains for later. However, in the mixed long multichains of HO-HE because of the presence of HO chains the dimerization is suppressed, the oscillation is enhanced for HO-HE ST and HO-HE TT. For the two configurations the oscillations with $N$ are consistent for odd $L$ and are opposite for even $L$ each other, such as when $L=3$ and $N=16$ $\eta$ is trough for both HO-HE TT and HO-HE ST, while $L=4$ and $N=16$ $\eta$ is peak for HO-HE ST and is trough for HO-HE TT. When the cavity is considered the oscillation is significantly strengthened for the configurations containing HE chain, as shown in Fig. 12(e-h). For HE ST the oscillation decreases with $L$, which is caused by the dissipation of the more and more TLSs. Whatever situation the oscillation of $\eta$ with $N$ for HE ST and HE TT is consistent, which illustrates further that the oscillation is mainly caused by HE chain. Interestingly, the oscillation amplitude behaves inverse dependence on the parity of $L$ between HE TT and HO-HE ST, HO-HE TT, i.e. for the former it with odd $L$ is bigger than that with even $L$, and for the latter it with even $L$ is bigger than that with odd $L$. Furthermore, the dependence of the exciton transport efficiency for HO-HE ST and HO-HE TT on the parity of $N$ can be changed by $L$ and the cavity. For example, comparing $L=2$ and $L=3$ for HO-HE ST of $N=16$, $\eta$ corresponds respectively to peak and trough [Fig. 12(a, b)], as well as comparing Fig. 12(b) and 12(f) for HO-HE ST of $N=14$ and $L=2$ it corresponds respectively to trough and peak. We also note by comparing Fig. 12(b) and 12(f), Fig. 12(f) and 12(g) for HO-HE ST, that the period of the oscillation with $N$ also is varied by the cavity and $L$.
V Summary
In summary, in order to reveal the mechanism of exciton transport in a 2D multichains system with a cavity, the spectra, the second order partial derivative of the Hopfield coefficients of cavity photons with respect to the inter-chain coupling strength and the dimerization parameter, the von Neumann entropy and the steady-state dynamics of exciton are obtained by numerically diagonalizing the Hamiltonian and solving the Lindblad quantum master equation, in which six inter-TLS coupling configurations are chosen as HO ST, HO TT, HE ST, HE TT, HO-HE ST and HO-HE TT based on the HO and HE chain, as well as the square and triangle type of inter-chain coupling. The effects of the inter-chain coupling strength, the dimerization parameter, the detuning of the driving frequency to the TLS, the number of chain, the length of chain, and the coupling configurations on them are investigated. The results show that in the absence of cavity the exciton transport efficiency is decided by the distribution of population of exciton on the whole multichains. Besides, when the cavity is considered the occupation number of the cavity photons depend significantly on the cavity-dressed energy level crossings and anticrossings of the nearly-zero-energy modes, at which the coherent excitation and LZ transition between the multichains and the cavity occur. Therefore, the exciton transport efficiency and currents via the polariton channel are changed at the locations of the crossings and anticrossings. We can conclude that the exciton transport dynamics are controlled by the crossings and anticrossings. Besides, the locations of the crossings and anticrossings is accurately determined by calculating von Neumann entropy. We also discover that the second order partial derivative of the Hopfield coefficients of cavity photons is a good physical quantity to describe the locations of the crossings and anticrossings, and the change of photonic occupation number. The spectra show that the locations of the crossings and anticrossings are related closely to the coupling configurations, the inter-chain coupling strength, and the dimerization parameter. Particularly, the exciton transport efficiency and currents are symmetric about the inter-chain coupling strength $\zeta=0$ in the multi-chain system without cavity, but the symmetry can be broken by the cavity. In the space of the dimerization parameter the presence of the cavity enhances exciton transport for HE ST, HO-HE ST and HO-HE TT with even chains, but suppresses it for HO-HE ST and HO-HE TT with odd chains when $\delta>0.36$. Interestingly, when the cavity drive is simultaneously exerted, the exciton transport is not limited by the occupation number of site $11$ and is strengthened for HE ST and HO-HE ST with odd chains in the range of the negative dimerization parameter. Further, in the space of the TLS transition frequency detuning and the inter-chain coupling strength we discover that the exciton transport is suppressed for HO ST and HO TT and is strengthened for the other configurations by the cavity at the range corresponding to the crossings and anticrossings, in which the polaritons act as vital role and the effects are enhanced by the cavity drive. In the space of the TLS transition frequency detuning and the dimerization parameter, the steady-state dynamics depend on the coupling configurations, the driven methods, and the cavity, especially in the zone of the fully dimerized limit. Finally, it can be found that the exciton transport is obviously related to the parity of the number and length of the chains with cavity, in which the exciton transport efficiency behaves oscillation with them. The oscillation is related to the configurations. In a word, our results reveal the exciton transport mechanism of the 2D hybrid system, which provides theoretical foundation for designing controllable and high efficient exciton transport devices, such as solar cell. Of course, it has many points of view to reveal the mechanism, like the topological phase transition of the system.
VI Acknowledgements
This work is supported by the Academic Ability Promotion Foundation for Young Scholars of Northwest Normal University in China under the Grant No. NWNU-LKQN2024-10, National Natural Science Foundation of China Regional Fund under Grant No. 12164042, National Natural Science Foundation of China under Grant No. 12104374, National Natural Science Foundation of China under Grant No. 12065022.