### v4

# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

Abstract
The performance of various quantum devices is fundamentally linked to the control of exciton transport. To explore this, we study the exciton transport of the two-dimensional multi-chain systems with different coupling configurations in an optical cavity. Two types of the chains—the homogeneous and heterogeneous coupling chain, as well as two inter-chain coupling conformations—the square and triangular arrangements, are considered. The effects of the inter-chain coupling, the dimerization parameter, the cavity, the length and number of the chains on exciton transport are systematically investigated for different coupling configurations through the spectra, the Hopfield coefficients, and the steady-state dynamics of the system. The results show that in the absence of a cavity the exciton transport currents and efficiency are determined by the exciton distribution across the multi-chain system. However, when a cavity is introduced the exciton transport can be significantly enhanced or suppressed by the polariton formation at the cavity-dressed energy level crossings and anticrossings near zero-energy modes, where the coherent excitation and Landau-Zener transitions occur. Meanwhile, we discover that the discontinuous and extremal points in the second-order partial derivatives of the photon Hopfield coefficients with respect to the inter-chain coupling and the dimerization parameter correspond respectively to the crossings and anticrossings at the extreme points of the photon occupation number. Additionally, exciton transport efficiency is closely related to the odevity of both chain length and chain number, and exhibits oscillatory behaviour. This work provides critical insights into the exciton transport mechanism in multi-chain–cavity system and theoretical basis for designing high-performance excitonic devices with tunable transport properties.
I Introduction
In quantum device, energy and information are primarily transported by three carriers—electron, photon, and exciton. Optimizing and controlling the efficiency, the speed, the distance, and the current of carrier transport are the center to realize the functions of various quantum devices, such as organic solar cells wurfel2015impact ; lu2022simultaneously ; melianas2017charge , organic light-emitting diodes shan2011transport ; tanase2003unification ; shen2022through , quantum dot solids zhang2022ultrafast , organic semiconductor devices sneyd2021efficient ; hagenmuller2017cavity , quantum simulation, and quantum computation georgescu2014quantum ; yang2019quantum etc. Exciton transport is influenced by multiple factors, including the dissipation cainelli2021exciton ; zhu2024quantum ; bose2022effect , exciton-exciton annihilation wittmann2020energy , disorder zhu2022single ; quenzel2022plasmon ; schachenmayer2015cavity ; allard2022disorder ; balasubrahmaniyam2023enhanced ; ribeiro2022multimode , defect sites wagner2011bound ; akselrod2014visualization , and topological property buendia2024long ; allard2023multiple ; nie2021topology . Consequently, extensive research focuses on revealing the exciton transport mechanisms and properties in diverse quantum systems engel2007evidence ; menke2013tailored , and improving transport properties by reducing and avoiding detrimental factors.
In experiment, the one-dimensional (1D) system carrying electron, photon, or exciton transport have been realized through various platforms, including organic aggregates molecule chainssu1979solitons , superconducting qubit circuits roushan2017spectroscopic ; xu2018emulating ; blais2003tunable , Rydberg-atom chains barredo2015coherent ; orioli2018relaxation , plasmons su2021perovskite ; seitz2020exciton , semiconductorssneyd2021efficient ; giannini2022exciton ; muller2023directed . At room temperature the strong exciton-photon coupling between organic aggregates and microcavities lidzey1998strong ; lidzey1999room ; kena2008strong ; tischler2005strong ; coles2014polariton or plasmons bellessa2014strong ; bellessa2004strong ; zengin2015realizing ; li2018plasmon can result in new applications like polariton lasing kena2010room , in which the polaritons form kavokin2003cavity ; agranovich2003cavity , and the molecular energy structures are modified galego2015cavity . The polariton formation significantly modulates exciton transport properties containing the efficiency wei2019enhanced ; wei2022cavity ; feist2015extraordinary , rate wang2021polariton ; aroeira2023theoretical , and distance du2018theory ; myers2018polariton ; sandik2024cavity ; takazawa2010fraction because the decay of excitons on chains is suppressed.
The hybrid system composed of a 1D molecular chain coupled to an optical cavity is well-described by the Tavis-Cummings model tavis1968exact . The chain with the staggered hopping interaction is depicted by the Su-Schrieffer-Heeger (SSH) model su1979solitons , which exhibits topologically trivial and nontrivial phases for even-numbered the two-level systems (TLSs) cai2019observation . In the topologically nontrivial phase, even-length chains host a small gap bewteen two localized nearly-zero-energy modes within the topological bandgap, enabling hybrid edge modes formation from the left and the right edge states. These topologically protected edge states enhance exciton transport in finite SSH chains, as confirmed theoretically and experimentally chang2020probing ; wei2022cavity . The gap’s origin lies in energy-level anticrossings, where the Landau–Zener (LZ) transitions occur. In the two-dimensional (2D) and three-dimensional tungsten disulfide material, the anti-crossing dispersions were presented, in which the enhancement of exciton transport capability was corroborated in experiment guo2022boosting ; yuan2017exciton . Theoretical study have explored cavity-coupled system with multichains kvande2023finite , suggesting that the exciton transport properties of the higher-dimensional ($\geq 1$) systems warrant further investigation.
Motivated by the recent studies on the high dimensional tungsten disulfide, in this paper we devote to the studies of the inter-chain joint effect on exciton transport in 2D TLSs array within a cavity. Each chain contains two alternative coupling types, i.e., the homogeneous (HO) and the heterogeneous (HE) coupling, and two conformations of the inter-chain coupling are set, i.e., square and triangle types. Firstly, the energy spectra of the strong exciton-cavity coupled system are obtained by diagonalizing numerically the Hamiltonian for six configurations, analyzing their dependence on inter-chain coupling strength and dimerization parameter. Meanwhile, we calculate the von Neumann entropy of exciton to confirm the locations of the crossings and anticrossings, and the second order partial derivative of the photon Hopfield coefficients to show the change of the occupation number of the cavity photon. What’s more, the effects of the inter-chain coupling, the dimerization parameter, the cavity, the TLS transition frequency detuning, the number and length of the chains on the steady-state dynamics of exciton transport are investigated for different configurations by numerically solving the Lindblad quantum master equation under the TLS and TLS+cavity drive. In the absence of cavity, the channel of exciton transport only is the multichains, in which its current and efficiency depend on the distribution of exciton on the whole multichains. Importantly, the exciton transport in the hybrid system correlates strongly with the crossings/anticrossings of dressed near-zero-energy mode, where the coherent excitation and LZ transition mediate the excitation exchange between the multichains and the cavity. Notably, polaritons play pivotal roles in controlling exciton transport through the crossings/anticrossings. In addition, the second order partial derivatives of the photon Hopfield coefficients serve as reliable metrics for identifying the extremal points in photon occupation number and the precise locations of the crossings/anticrossings. Further, we present the optimal parameters regions with the TLS transition frequency detuning, the inter-chain coupling strength, and the dimerization parameter. Finally, the exciton transport dynamics exhibit the odevity-dependent oscillations tied to the number and length of the chains.
The remainder of this paper is organised as follows: Sec. II presents the theoretical framework, including the Hamiltonian of the multichain-cavity coupling system (in Sec. II.1), Von Neumann entropy formulation for excitons (in Sec. II.2), the Lindblad quantum master equation of the system with an external driving field (in Sec. II.3), and the steady-state exciton transport currents and efficiency (in Sec. II.4). In Sec. III the energy spectra are given by numerically diagonalizing the Hamiltonian. In Sec. IV we simulate the steady-state exciton dynamics across multi-chain configurations, analyzing the effects of the cavity, the inter-chain coupling, the dimerization parameter, the TLS transition frequency detuning (in Sec. IV.1), the number and length of the chains (in Sec. IV.2) on exciton transport. Section V is a brief summary and future perspective.
Figure 1: (color online) The schematic representation of the hybrid system consisting of the 2D TLSs array within a cavity for three configurations HO ST (a), HO TT (b), and HE ST (c). Red spheres represent the TLSs with ground state $|g\rangle$ and excited state $|e\rangle$ between which the transition frequency is $\omega_{\rm a}$ and the decay rate is $\gamma$. The decay rate of the imperfect cavity is $\kappa$. The 2D TLSs array can be driven by an external coherent field with driving frequency $\omega$ via the first TLS of the first row with driving amplitude $\xi_{{\rm T},1}$ or/and via the cavity with driving amplitude $\xi_{c}$. The excitons in the multichains finally outflow from the last column into a reservoir (drain) with rate $\gamma_{\rm R}$. $h$ is the intra-chain hopping strength between adjacent TLSs. $\delta$ is the dimerization strength of HE chain. $\zeta$ is the inter-chains couplings strength between the nearest TLSs.
II Model
II.1 Hamiltonian
The considered system is a 2D molecular aggregate embedded in a single-mode cavity, which can be modelled as multiple parallel chains with number $L$ ($L\geqslant 2$) as illustrated in Fig. 1. The $k$th chain comprises $N_{k}$ identical TLSs. For simplicity, we assume that all chains contain the same number of the TLSs ($N_{1}=N_{2}=...=N_{L}=N$). The Hamiltonian of the multichain-cavity coupled system can be expressed as
$H_{0}=H_{\rm e}+H_{\rm c}+H_{\rm ec}$ (1)
The $H_{\rm e}$ represents the Hamiltonian of the multichains, and contains two terms as
$H_{\rm e}=\sum^{L}_{k=1}H_{{\rm e},k}+H_{\rm e,inter-chain},$ (2)
where $H_{{\rm e},k}$ are the term of the $k$th bare chain.
Considering the staggered hopping interaction between the nearest-neighbor TLS monomers in a chain, under the rotating frame of the driving frequency $\omega$ the Hamiltonian of the $k$th chain is described by the SSH model
$H_{{\rm e},k}=$ $-\sum_{i=1}^{N}\Delta_{\rm a}\sigma_{k,i}^{+}\sigma_{k,i}^{-}-$
$\sum_{i=1}^{N-1}h\left[1+(-1)^{i}\delta\right](\sigma_{k,i}^{+}\sigma_{k,i+1}^{-}+{\rm H.c.}),$ (3)
where $h$ characterizes the hopping amplitude between two nearest-neighbor TLS monomers in a chain, $\delta$ represents the dimerization strength, $\sigma_{k,i}^{+}=|e_{k,i}\rangle\langle g_{k,i}|$ ($\sigma_{k,i}^{-}=|g_{k,i}\rangle\langle e_{k,i}|$) is the raising (lowering) operator of the TLS with same energy $\omega_{\rm a}$ (setting $\hbar=1$) at the $i$th site of the $k$th chain with the ground state $|g_{k,i}\rangle$ and the excited state $|e_{k,i}\rangle$ because of the identical particle, $\Delta_{\rm a}=\omega-\omega_{\rm a}$ is the TLSs transition frequency detuning relative to driving field frequency $\omega$, and ${\rm H.c.}$ denotes the Hermitian conjugate. The SSH chain behaves entirely different property for $\delta=0$ and $\delta\neq 0$ corresponding to homogeneous and heterogeneous coupling in a molecular chain respectively, because the topologically nontrivial regime exists for the latter. In order to emphasize the difference in exciton transport between the two coupling cases, we designate separately them as the homogeneous (HO) and the heterogeneous (HE) coupling chain in the following content.
The inter-chain coupling term is denoted by the $H_{\rm e,inter-chain}$, whose form depends on the inter-chain coupling configuration in the 2D multichains. In this paper, we consider the two simplest configurations, the square type (ST) [Fig. 1(a)] and the triangle type (TT) [Fig. 1(b,c)]. The corresponding Hamiltonian can be respectively written as
$H_{\rm e,inter-chain}=$ $\zeta\sum^{L-1}_{k=1}\sum_{i=1}^{N}(\sigma_{k,i}^{+}\sigma_{k+1,i}^{-}+{\rm H.c.})$ for ST, (4a)
$H_{\rm e,inter-chain}=$ $\zeta\sum^{P}_{k=1}\sum_{i=1}^{N-1}[\sigma_{2k-1,i}^{+}(\sigma_{2k,i}^{-}+\sigma_{2k,i+1}^{-})+\sigma_{2k-1,N}^{+}\sigma_{2k,N}^{-}+{\rm H.c.}]+$
$H_{\rm e,inter-chain}=$ $\zeta\sum^{Q}_{k=1}\sum_{i=2}^{N}[\sigma_{2k,1}^{+}\sigma_{2k+1,1}^{-}+\sigma_{2k,i}^{+}(\sigma_{2k+1,i-1}^{-}+\sigma_{2k+1,i}^{-})+{\rm H.c.}]$ for TT, (4b)
where for TT the upper bounds of $L$ in odd and even term are respectively
$P=$ $(L-1)/2$ $L$ is odd, (5a)
$P=$ $L/2$ $L$ is even, (5b)
$Q=$ $(L-1)/2$ $L$ is odd, (6a)
$Q=$ $(L-2)/2$ $L$ is even, (6b)
and $\zeta$ denotes the inter-chain coupling strength, namely the hopping amplitude between two nearest-neighbor TLSs in two nearest-neighbor chains. Therefore, the multi-chain system exhibits six possible configurations based on the permutation and combination of the inter-chain and the intra-chain coupling types, which are expressed as HO ST, HO TT, HE ST, HE TT, HOCC-HE ST, and HO-HE TT. The HO ST, HO TT, and HE ST are schematically depicted in Fig. 1(a), Fig. 1(b), and Fig. 1(c) respectively. The HE TT configuration is obtained by replacing the HO chains in HO-TT with HE chains. For the hybrid cases, the HO and HE chains are arranged in a staggered pattern in both HO-HE ST and HO-HE TT, in which the odd-numbered and even-numbered chains are respectively HO and HE type.
In the rotating frame of the driving frequency $\omega$, the cavity radiation field is described by the Hamiltonian
$H_{\rm c}=-\Delta_{\rm c}c^{\dagger}c,$ (7)
where $c^{\dagger}=|c\rangle\langle 0|$ ($c=|0\rangle\langle c|$) represents the photon creation (annihilation) operator, $\omega_{\rm c}$ is the cavity-mode frequency, $\Delta_{\rm c}=\omega-\omega_{\rm c}$ is the detuning of $\omega_{\rm c}$ with respect to $\omega$, $|c\rangle$ denotes the single-photon excited state of the cavity field, and $|0\rangle$ represents the joint vacuum state of the exciton-cavity system. To ensure the resonant condition between the cavity and the TLSs, we set $\omega_{\rm c}=\omega_{\rm a}$, hence $\Delta_{\rm a}=\Delta_{\rm c}=\Delta$.
In the rotating frame of $\omega$, the light-matter coupling between the multi-chain aggregate and the cavity is given by
$H_{\rm ec}=\sum^{L}_{k=1}\sum_{i=1}^{N}g_{k,i}(\sigma_{k,i}^{+}c+{\rm H.c.}),$ (8)
where $g_{k,i}$ represents the coupling strength between the TLS at the $i$th site in the $k$th chain and the cavity field. For the sake of simplicity, we assume uniform coupling $g_{k,i}=g$ for all TLSs. Generally, $g$ depends on the collective Rabi splitting $\Omega_{\rm R}=2\sqrt{\sum_{k=1}^{L}\sum_{i=1}^{N}g^{2}_{k,i}}=2g\sqrt{LN}$.
In order to investigate the excitation dynamics, we introduce an external coherent drive in the exciton-cavity system. We consider two driving means, i.e., the driving field is exerted on the first site of the $k$th chain, as well as simultaneously on both this site and cavity, which can be described respectively by
$H_{\rm d}=$ $H_{\rm dT}=\xi_{{\rm T},k}(\sigma_{k,1}^{-}e^{i\omega t}+\sigma_{k,1}^{+}e^{-i\omega t}),$ (9a)
$H_{\rm d}=$ $H_{\rm dTc}=H_{\rm dT}+\xi_{\rm c}(ce^{i\omega t}+c^{\dagger}e^{-i\omega t}),$ (9b)
where $\omega$ is the driving frequency, $\xi_{{\rm T},k}$ ($\xi_{\rm c}$) denotes the driving strength for the TLS (cavity). The second mean is denoted by ”TLS+cavity” in the following. Here, we set $k=1$, namely, only the first site in the first chain is driven. The two driving means were used to study the two-photon blockade in the cavity QED system hamsen2017two . In the rotating frame of $\omega$, the Eq. (9) can be written as
$H_{\rm d}=$ $H_{\rm dT}=\xi_{{\rm T},1}(\sigma_{1,1}^{-}+\sigma_{1,1}^{+}),$ (10a)
$H_{\rm d}=$ $H_{\rm dTc}=H_{\rm dT}+\xi_{c}(c+c^{\dagger}).$ (10b)
When the driving field is considered, the total Hamiltonian of the open quantum system is
$H=H_{0}+H_{\rm d}.$ (11)
[…]
II.2 von Neumann entropy
In order to find and distinguish accurately the locations of the energy level crossings and anticrossings of Hamiltonian (1) in the parameter space of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$, one can calculate the von Neumann entropy of the excitons by
$S(\rho_{\rm e})=-{\rm tr}(\rho_{\rm e}{\rm log}_{2}\rho_{\rm e}).$ (12)
For the exciton-photon hybrid system, its eigenstates can be indicated as $|\psi_{\rm e-p}\rangle$. Thus the reduced density matrix $\rho_{\rm e}$ of the excitons is obtained by tracing out the degrees of freedom of the photons, i.e., $\rho_{\rm e}={\rm tr}_{\rm p}(|\psi_{\rm e-p}\rangle\langle\psi_{\rm e-p}|)$. According to Ref. hu2018effect , the crossings (anticrossings) correspond to the saltatorial intersections (maxima or acclive intersections) of the von Neumann entropy with parameter. Therefore, the entanglement between upper and lower energy levels is maximal at the anticrossings, where the LZ transitions occur. The transition from anticrossing to crossing means the energy degeneracy hu2018effect . At the crossings, the excitations are exchanged coherently between multichains and cavity. As a result, the boosted exciton transport in degenerate states was demonstrated tagarelli2023electrical ; lin2012efficient .
II.3 Transport dynamics
For the system that exists dissipation at each TLS in the multichains, the driving field, and the coupling with environment (reservoir), its dynamics are governed by the Lindblad quantum master equation under the Born-Markov approximations and the rotating-wave approximation lindblad1976generators . Then, the exciton transport dynamics for the Hamiltonian (11) can be investigated by the master equation
$\dot{\rho}=i[\rho,H]+\sum^{L}_{k=1}(D_{{\rm d},k}[\rho(t)]+D_{{\rm R},k}[\rho(t)]),$ (13)
where the first term in the righthand side accounts for the coherent evolution under the total Hamiltonian (11), the second term characterizes the decay of excitons due to the spontaneous emission with
$D_{{\rm d},k}[\rho]=\sum^{N}_{i=1}\gamma_{d}\mathcal{L}_{\sigma^{-}_{k,i}}[\rho]+\kappa\mathcal{L}_{\rm c}[\rho],$ (14)
where $\gamma_{d}$ is the spontaneous emission rate of the TLSs, $\kappa$ is the decay rates of the cavity photons, and $\mathcal{L}_{x}[\rho]=x\rho x^{\dagger}-\dfrac{1}{2}\left\{\rho,x^{\dagger}x\right\}$ is the Lindblad superoperators with standard form. The term $D_{{\rm R},k}[\rho(t)]$ in Eq. (13) describes the exciton exchange between chain ends and exciton drain, which is given by
$D_{{\rm R},k}[\rho]=\gamma_{\rm R}\bar{n}_{\rm R}\mathcal{L}_{\sigma_{k,N}^{+}}[\rho]+\gamma_{\rm R}(\bar{n}_{\rm R}+1)\mathcal{L}_{\sigma_{k,N}^{-}}[\rho],$ (15)
where $\gamma_{\rm R}$ is the exciton exchange rate, $\bar{n}_{\rm R}$ is the average exciton number in the drain. The last sites of each chain exist coupling with the drain, so $k\in\{1,2,...,L\}$. The two terms in Eq. (15) represent respectively the inflow and outflow of excitons from the last column TLSs of the multichains to the drain. In order to show the net transportation of exciton, we set $\bar{n}_{\rm R}=0$, establishing unidirectional exciton transfer from the last column of the multichains to the drain without reflection, which is reasonable as the photosynthesis process in nature.
II.4 Efficiency of exciton transport
Figure 2: (color online) The energy spectra of the exciton-cavity coupling system as a function of the inter-chain coupling strength $\zeta$ for different configurations HO ST (a), HE ST (b), HO-HE ST (c), with the colormap representing the photon Hopfield coefficients by numerically diagonalizing Hamiltonian (1) in single-excitation space. (d-f) The second order partial derivative of the photon Hopfield coefficients with respect to $\zeta$, which contains the locations of discontinuity and extremum corresponding to that existing evident changes in exciton transport dynamics. (g-i) The corresponding von Neumann entropy $S$ of excitons by calculating Eq. (12), in which the colors of the lines has not means just to distinguish each other. The black (violet) vertical lines and the data texts mark the locations of cavity-dressed energy level crossings (anticrossings) (a-c) producing obvious effect on exciton transport, the discontinuities (extrema) of the second order partial derivative of the photon Hopfield coefficients (d-f), as well as the saltatorial intersections (maxima or acclive intersections) of the von Neumann entropy (g-i). The corresponding parameters are $L=2$, $N=8$, $\Delta_{\rm a}=\Delta_{\rm c}=0$ eV, $h=1$ eV, $\delta=0.5$, and $\Omega_{R}=1$ eV.
Figure 3: (color online) Similar as Fig 2, but for different configurations HO TT (a, d, and g), HE TT (b, e, and h), and HO-HE TT (c, f, and i).
Figure 4: (color online) Similar as Fig 2, but for even chain $N=8$ and for different configurations HE ST (a, e, and i), HE TT (b, f, and j), HO-HE ST (c, g, and k), and HO-HE TT (d, h, and l) with dimerization parameter $\delta$. The parameter used is $\zeta=0.1$ eV, the others are the same as those in Fig. 2.
Figure 5: (color online) Similar as Fig 4, but for odd chain $N=9$.
To quantify the exciton transport dynamics in the 2D multichain-cavity system, the exciton transport efficiency $\eta$ is defined and derived in this subsection. Generally, the exciton transport efficiency is characterized by the ratio of the outgoing $(I^{\rm s}_{\rm o})$ and input $(I^{\rm s}_{\rm i})$ exciton currents for steady state liu2019vibration ; wei2019enhanced , that is
$\eta=\lim_{t\to\infty}\frac{I_{\rm o}}{I_{\rm i}}=\frac{I_{\rm o}^{\rm s}}{I_{\rm i}^{\rm s}},$ (16)
where the superscript ”s” refers to the steady state. On the other hand, the exciton occupation number can reveal the transport mechanism and determine the currents. The exciton occupation number on the $j$th site of the $k$th chain is given by
$p_{k,j}(t)\equiv{\rm Tr}\{\rho\sigma^{+}_{k,j}\sigma^{-}_{k,j}\}.$ (17)
Combining Eq. (13) and Eq. (17), the time evolution of $p_{k,j}(t)$ obeys
$\dot{p}_{k,j}(t)=$ $i{\rm Tr}\{[\rho,H]\sigma^{+}_{k,j}\sigma^{-}_{k,j}\}+$
${\rm Tr}\{\sum_{\mu={\rm d,R}}\sum^{L}_{k=1}D_{\mu,k}[\rho]\sigma^{+}_{k,j}\sigma^{-}_{k,j}\},$ (18)
where the trace is operated throughout the Hilbert space of the total system. The lefthand side describes the rate of the change of $p_{k,j}(t)$ with time, thus the first term (the second term) of the righthand side represents the input (outgoing) exciton current at the $j$th site in the $k$th chain. According to the explanation, under the TLS and TLS+cavity drive the input exciton current of the multichains can be written respectively as
$I_{\rm i}(t)=$ $i{\rm Tr}\{[\rho,H_{\rm dT}]\sigma^{+}_{1,1}\sigma^{-}_{1,1}\},$ (19a)
$I_{\rm i}(t)=$ $i{\rm Tr}\{[\rho,H_{\rm dT}]\sigma^{+}_{1,1}\sigma^{-}_{1,1}+[\rho,\xi_{c}(c+c^{\dagger})]c^{\dagger}c\},$ (19b)
which is related to the exciton injection at the first site in the $k$th chain or/and in cavity via the external coherent field per unit time. Substituting Eq. (9) for $H_{\rm dc}$ and $H_{\rm dTc}$ in Eq. (19), under the zero- and single- excitation subspace, the input exciton currents are simplified to
$I_{\rm i}(t)=$ $i\xi_{{\rm T},1}(\langle 1,1|\rho|0\rangle-{\rm H.c.}),$ (20a)
$I_{\rm i}(t)=$ $i\xi_{{\rm T},1}(\langle 1,1|\rho|0\rangle-{\rm H.c.})+i\xi_{\rm c}(\langle c|\rho|0\rangle-{\rm H.c.}).$ (20b)
According to the Hermitian property of density matrix, under the TLS and TLS+cavity drive the input exciton currents are finally written as
$I_{\rm i}(t)=$ $2\xi_{{\rm T},1}{\rm Im}[\langle 0|\rho|1,1\rangle],$ (21a)
$I_{\rm i}(t)=$ $2\xi_{{\rm T},1}{\rm Im}[\langle 0|\rho|1,1\rangle]+2\xi_{c}{\rm Im}[\langle 0|\rho|c\rangle].$ (21b)
Analogous to the input exciton current, the outgoing exciton current can be defined as the outflow of exciton transfering from the last column in the multichains to the drain per unit time. According to the means of the third term of righthand in Eq. (18), it can be given by
$I_{\rm o}(t)={\rm Tr}[\sum^{L}_{k=1}D_{{\rm R},k}[\rho(t)]\sigma_{k,N}^{+}\sigma_{k,N}^{-}].$ (22)
Pluging the Eq. (15) into the Eq. (22), under the single-excitation assumption, the form of the outgoing exciton current is simplified as
$I_{\rm o}(t)=\gamma_{\rm R}\sum^{L}_{k=1}p_{k,N},$ (23)
where $p_{k,N}$ is the occupation number at the last site of the chain $k$.
III Spectrum of coupling system
There are a large number of studies on the spectra of the SSH Hamiltonian for the open boundary conditions sirker2014boundary ; nie2020bandgap ; wei2022cavity ; cai2019observation . The topological phases of the SSH chain depend on the odevity of TLS number in 1D chain and the qubit-qubit coupling configuration. According to the Ref. cai2019observation , the dimerization parameter $\delta>0$ ($\delta<0$) corresponds to the topological winding number $\nu=1$ ($\nu=0$), i.e., the topologically nontrivial (trivial) regime. In Ref. sirker2014boundary , the spectra show that there is single zero-energy mode in whole range of $\delta$ for odd-length chain and are two localized nearly-zero-energy modes in the topologically nontrivial regime $\delta>0$ for even-length chain. The finite gap between the two edge modes leads to the formation of the hybrid edge modes that overlap with each other, generating the polaritonic states when the edge modes are dressed by cavity due to the exciton-photon coupling. Comparing to the single chain system, the presence of the inter-chain coupling induces the spectral extension of the 2D multi-chain system, adding greatly to the complexity of the system dynamics. In the exciton-cavity coupling system, the crossings and anticrossings of the edge modes dressed by the cavity can lead to the abundant exchange of excitation between the multichains and the cavity attributing to the coherent exchanges and the LZ transitions, where the polaritons increase or decrease. The polaritons act as intermediary in exciton transport, causing that the excitons are transported among site $(k,1)~{}\rightarrow$ cavity $\rightarrow$ site $(k,N)$ via bypassing dissipative intermediate TLSs wei2022cavity . Therefore, the exciton transport can be enhanced by the polaritons at the crossings and anticrossings.
Figure 6: (color online) The dependence of the steady-state occupation numbers (a1 and b1) $p_{11}$ on site (1,1), (a2 and b2) $p_{\rm bulk}$ on bulk sites, (a3 and b3) input exciton current $I_{\rm i}$, (a4 and b4) outgoing exciton current $I_{\rm o}$, and (a5 and b5) transport efficiency $\eta$ on the inter-chain coupling strength $\zeta$ (a1-5) and the dimerization parameter $\delta$ (b1-5) for different configurations without cavity. In (b1-5) the solid and dash lines correspond respectively to $N=8$ and 9. The parameters used are $\zeta=0.1$ eV, $\xi_{{\rm T},1}=\xi_{\rm c}=0.1$ eV, $\gamma_{\rm d}=0.01$ eV, $\gamma_{\rm R}=0.1$ eV, and the others are the same as those in Fig. 2.
Figure 7: (color online) The dependence of the steady-state dynamics of the hybrid system on the inter-chain coupling strength $\zeta$ in the presence of cavity under TLS drive (a1-6) and TLS+cavity drive (b1-6). (a5 and b5) The occupation number of photon with $\zeta$, in which the extreme points are marked by the vertical fine lines, and the others are similar to Fig. 6. The parameter used is $\kappa=0.01$ eV, and the others are the same as those in Fig. 6.
Figure 8: (color online) The dependence of the steady-state dynamics of the hybrid system on the dimerization parameter $\delta$, and the others are similar to Fig. 7. The parameters used are the same as those in Fig. 7.
In order to analyse the spectral characteristics of the multichain-cavity coupled system, we numerically diagonalize the Hamiltonian (1) to obtain its eigenvalues. The formation of polaritons is investigated through the photon Hopfield coefficients $|\alpha_{c}|^{2}$. When the $j$th normalized eigenvector of the Hamiltonian (1) is expanded as $|\psi_{j}\rangle=\Sigma_{k}^{L}\Sigma_{i}^{N}\alpha_{k,i}|k,i\rangle+\alpha_{c}|c\rangle$, $|\alpha_{k,i}|^{2}$ and $|\alpha_{c}|^{2}$ denote the Hopfield coefficient of exciton and photon, which characterize the hybridization weights of exciton and photon components, respectively. The color maps Fig. 2 (a-c), 3 (a-c) and Fig. 4 (a-d), 5 (a-d) show respectively the spectra of the exciton-cavity coupled system as functions of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$ for different coupling configurations, and contain the photon Hopfield coefficient $|\alpha_{c}|^{2}$. One can see that the zero-energy modes exist in the spectra for all situations. $|\alpha_{c}|^{2}\neq 0$ means that the energy level is dressed by the cavity mode. It is obvious that only the nearly-zero-energy modes are distinctly dressed by the cavity mode, generating corresponding polaritonic branches, while the high excited states are hardly dressed for the given exciton-cavity coupling strength. In addition, there are the crossings (anticrossings) in the spectra, in which the crossings (anticrossings) between the nearly-zero-energy modes dressed by the cavity mode are our focus. It can be seen that the photonic content distribution on the upper and lower polaritonic branches varies significantly with $\zeta$ and $\delta$. It means that the coherent exchange of excitation or the LZ transition of photons occurs, inducing the changes of the exciton transport efficiency and currents, which will be analysed in the next section by the steady-state dynamics. Further, it can be discovered that the number and locations of crossings (anticrossings) affecting transport depend on the coupling configurations of TLSs in the range of $\zeta$ and $\delta$. In comparison, the effect of $\zeta$ on the crossings (anticrossings) is more significant than that of $\delta$.
To systematically analyze the photon Hopfield coefficient $|\alpha_{c}|^{2}$, we plot the scatter maps of the second order partial derivatives $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in Fig. 2 (d-f) and 3 (d-f), as well as $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ in Fig. 4 (e-i) and 5 (e-i), which correlate with the obvious changes of the steady-state exciton transport dynamics. To determine and distinguish accurately the locations of the crossings and anticrossings in spectra, the von Neumann entropy $S$ of excitons is calculated through Eq. (12), and is shown in Fig. 2 (g-i), 3 (g-i), 4 (i-l), and 5 (i-l). The dressed nearly-zero-energy crossings (anticrossings) identified by $S$ are marked in the spectra, which correspond to the saltatorial intersections (maxima or acclive intersections) of $S$. Significantly, we discover that the discontinuous locations of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with $\zeta$ and $\delta$ correspond to the dressed nearly-zero-energy crossings, and the extrema points of that correspond to the dressed nearly-zero-energy anticrossings. As shown in Fig. 2(a,d,g), the discontinuities of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ at $\zeta=-0.3$ in green and red lines align with the crossings, which correspond to the saltatorial intersections of $S$. The maxima of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in green (at $\zeta=-0.21$) and cyan (at $\zeta=-0.02$) lines, and the minima of that in blue (at $\zeta=-0.21$) and green (at $\zeta=-0.02$) lines appear at the anticrossings, which correspond to the maxima of $S$. Similar patterns emerge throughout our spectral results. We also can see that the discontinuities in green (at $\zeta=0.07$), blue (at $\zeta=0.09$), and cyan (at $\zeta=0.22$) lines [Fig. 2(e)] match the crossings [Fig. 2(b)] at the saltatorial intersections of $S$ [Fig. 2(h)]. The maxima in green (at $\zeta=-0.03$) and magenta (at $\zeta=0.48$) lines, and the minima in magenta (at $\zeta=-0.03$) and cyan (at $\zeta=0.48$) lines [Fig. 2(e)] indicate the anticrossings [Fig. 2(b)] at the maxima of $S$ [Fig. 2(h)]. In Fig. 2(c,f,i), the maxima of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in black (at $\zeta=-0.1$), green (at $\zeta=-0.05$), and blue (at $\zeta=0.03$) lines, and the minima of that in red (at $\zeta=-0.05$) and cyan (at $\zeta=0.03$) lines correspond to the anticrossings indicated by the maxima of $S$. Therefore, the second order partial derivatives of the photon Hopfield coefficients serve as robust indicators for identifying the spectral crossings and anticrossings. About $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ in Fig. 3 (d-f), and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ in Fig. 4 (e-i) and 5 (e-i), the conclusion is similar as that of Fig. 2, thus it is no longer stated hereinbelow.
Comparing the spectra under different configurations, one can discover that the gap in the range $\zeta\in[-0.21,-0.02]$ for the configuration HO ST and its range [Fig. 2(a)] are respectively widened and extended by the heterogeneous coupling in multichains as shown in Fig. 2(b). For the mixed configuration HO-HE ST the new gaps near $\zeta=0$ appear as shown in Fig. 2(c). When the inter-chain coupling is TT, the spectra are shown in Fig. 3(a-c). It can be seen that the gaps near $\zeta=0$ narrow, and their ranges in $\zeta$ reduce compared with Fig. 2(a-c). In the space of $\delta$, the spectra are presented in Fig. 4 (a-d) for even chain and in Fig. 5 (a-d) for odd chain. In Fig. 4 (a-d) we can see that the configuration HE ST has more crossings of the dressed nearly-zero-energy modes than HE TT. What’s more, the crossings at $\delta=0.3$ and 0.36 [Fig. 4 (c)] turn into anticrossing at $\delta=0.19$ [Fig. 4 (d)] because of TT coupling. In the mixed multichains, the presence of the HO chain leads also to the decrease of the crossings, comparing Fig. 4 (a) and 4 (c). Comparing Fig. 4 with Fig. 5, one can discover that the dressed nearly-zero-energy levels cross for even-length chain and avoid crossing for odd-length chain. Therefore, we can suppose that the coherent exchange is main mechanism affecting exciton transport for former, while the LZ transition is for latter.
In the next section, the steady-state dynamics with $\zeta$ and $\delta$ will be shown, in which the relation between the dressed nearly-zero-energy crossings (anticrossings) and the occupation number of photons are discussed. The exciton transport mechanism are revealed. The changes of exciton transport currents and efficiency with $\zeta$ and $\delta$ are explained. The optimal ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$ for exciton transport are found out by comparing the spectra and steady-state dynamics for different configurations.
Figure 9: (color online) The dependence of the steady-state dynamics on the inter-chain coupling strength $\zeta$ and the detuning $\Delta$ without cavity under TLS drive (a-f, ab-fb), and with cavity under TLS drive (g-l, gc-lc) and under TLS+cavity drive (m-r) for HO ST (the 1st column), HO TT (the 2ed column), $\cdots$, and HO-HE TT (the 6th column). The physical quantities of each row are written at right. The zones that the cavity induces significant effect on transport efficiency $\eta$ are marked by ellipses in (g-l), the corresponding zones are marked in (ab-fb) and (gc-lc). The parameters used are the same as those in Fig. 7.
Figure 10: (color online) The dependence of the steady-state transport efficiency $\eta$ on the dimerization parameter $\delta$ and the detuning $\Delta$, in which the conditions of the column and row are written at the top and right of this figure respectively. The parameters used are the same as those in Fig. 7.
Figure 11: (color online) The steady-state dynamics as a function of the chain number $L$ without (a1-g1) and with (a2-h2) cavity for different configurations, in which $p_{1,i}$, $p_{L,i}$, and $p_{\rm bulk-in}$ represent respectively the occupation number of the first row, the last row, and the bulk state denoted as ”bulk-in” containing all TLSs inside of the multichains. The parameter used is $N=10$, and the others are the same as those in Fig. 7.
Figure 12: (color online) The steady-state dynamics as a function of the chain length $N$ for different configurations and chain number under TLS drive, (a-d) without cavity, (e-h) with cavity. The parameters used are the same as those in Fig. 7.
IV The exciton transport dynamics by multichains
In this section, we discuss the effects of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, the number $N$ and length $L$ of the chains on the exciton transport dynamics by numerically solving the master equation Eq. (13) under an external driving, which are displayed in two subsections below. According to the analysis in the Sec. III, the nearly-zero-energy modes are dressed by the cavity for strong exciton-cavity coupling, generating the polaritons. To ensure the polariton-dominated exciton transport, we set that the Rabi splitting ($\Omega_{R,k}$) is far stronger than both the cavity ($\kappa$) and TLSs ($\gamma$) decay rates, i.e., $\Omega_{R,k}\gg(\kappa,\gamma)$.
IV.1 The effects of the inter-chain coupling and the dimerization parameter
Here, we investigate primarily the corresponding relationship between the cavity-dressed energy level crossings (anticrossings) and the exciton transport dynamics to elucidate the underlying transport mechanism. Based on the relation and the properties of exciton transport dynamics, we identify the optimal parameter ranges of the inter-chain coupling $\zeta$, the dimerization parameter $\delta$, and the TLS transition frequency detuning $\Delta$, as well as the optimal coupling configurations for enhanced exciton transport. Meanwhile, there are two physical quantities to be introduced, the steady-state occupation number $p_{11}$ of site (1,1) and $p_{\rm bulk}$ of bulk sites, where the site (1,1) denotes the first row and column in the TLS array shown in Fig. 1, and the ”bulk” site contains collective occupation of interior sites (columns 2 to N-1). This notation convention applies throughout unless specified otherwise. In addition, according to the Eq. (23) the outgoing exciton current $I_{\rm o}$ is directly proportional to the total occupation number $\Sigma_{1}^{L}p_{k,N}$ at terminal sites. Hence, subsequent figures display only $I_{\rm o}$ as the transport metric, and $\Sigma_{1}^{L}p_{k,N}$ is denoted by $p_{N}$.
Fig. 6 shows the dependence of the steady-state exciton transport dynamics of the multi-chain system without cavity on $\zeta$ and $\delta$ for the six chosen configurations. At first glance, one can discover that the inter-chain coupling and dimerization parameter exhibit distinct influences on exciton transport. The $p_{11}$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ with $\zeta$ are symmetric about $\zeta=0$ in $[-1,1]$, but with $\delta$ are not.
For the six configurations, the descending sort of efficiency is successively HO-HE ST, HO ST, HO TT, HO-HE TT, HE ST, HE TT in $\zeta\in[-1,1]$ [Fig. 6(a5)]. For the HO-HE ST, near $\zeta=0.3$ the sites of bulk [Fig. 6(a2)], input, and outgoing possess simultaneously relatively high occupation numbers of exciton, which support exciton transport in multichains. Consequently, this behaves the largest exciton transport efficiency, input and outgoing currents [Fig. 6(a3,4,5)]. On the contrary, at $\zeta=0$ and $\zeta\rightarrow\pm 1$, despite more excitations at the site (1,1) [Fig. 6(a1)], the exciton occupation number $p_{N}$ at outgoing sites is too few to transfer to the drain, leading to the low efficiency and exciton currents. Specially, the excitons are distributed mainly at bulk sites for $\zeta\rightarrow\pm 1$ [see cyan line in Fig. 6(a2)], which is beneficial to energy storage and has potential applications for designing quantum battery quach2022superabsorption . Except for the HE ST, the other configurations also have the property for $\zeta\rightarrow\pm 1$, as well as HO TT and HO-HE TT do near $\zeta=\pm 0.4$. However, for HO ST near $\zeta=\pm 0.35$ and ±$\pm 0.97$, $p_{\rm bulk}$, $I_{\rm i,o}$, and $\eta$ are relatively high, thus the abilities both of energy transport and storage are strong. Near $\zeta=\pm 0.7$ for HO ST, although the exciton transport efficiency is large, exciton currents are relatively small. Thus it is not beneficial to the exciton transport.
The system has four configurations HE ST, HE TT, HO-HE ST, HO-HE TT containing HE chain. Because of the different spectra between odd- and even-length HE chain, Fig. 6(b1-5) compares the steady-state dynamics with $\delta$ for $N=8,~{}9$ in the absence of cavity. It is obvious that the exciton currents and transport efficiency are the largest near $\delta=0.3$ for the configurations HE ST and HO-HE ST of $N=8$ than that in the other range of $\delta$ and for the other configurations [Fig. 6(b3-5)], which are good for effective exciton transport. However, the HE TT of $N=9$ is the most beneficial to energy storage because of the high $p_{\rm bulk}$ [Fig. 6(b2)] as well as the lowest $I_{\rm i,o}$ and $\eta$. Particularly, the exciton transport is anomalous in $\delta<0$, where the efficiency is high, but the currents approach to zero, as shown in Fig. 6(b3-5), thus there is no real exciton transport. The reason is that the drive fails to coherently excite the site (1,1), namely the excitons can not be injected in multichains [Fig. 6(b1,2)]. When $\delta>0.35$, the exciton currents and transport efficiency decrease sharply with $\delta$ or too small, because the excitons are localized at the left-edge site so that the occupation number of the right-edge sites is small for the HO-HE ST and HO-HE TT of $N=8,9$ or approaches to zero for the others.
Considering a cavity, we plot the Fig. 7 and 8 corresponding to Fig. 6, which present the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ on the exciton transport dynamics. Our analysis specifically focuses on the change of the steady-state dynamics with $\zeta$ and $\delta$ at the cavity-dressed nearly-zero-energy crossings and anticrossings. In the multichain-cavity coupled system, the main channels of exciton transport contain the multichains and polaritons which depend on occupation number of excitons and photons, while it only is the multichains in the absence of the cavity.
Under the TLS and the TLS+cavity drive, the effect of the inter-chain coupling $\zeta$ on the steady-state dynamics of the multi-chain system in a cavity is presented for different configurations in Fig. 7(a1-6) and 7(b1-6), respectively. Particularly, to analyse what role the cavity acts as in exciton transport, the dependence of the photon occupation number $p_{\rm c}$ on $\zeta$ is shown in Fig. 7(a5) and 7(b5), in which the extreme points of $p_{\rm c}$ are marked by the vertical fine lines. It can be found through comparison that the extreme points correspond to the dressed nearly-zero-energy crossings and anticrossings in spectra, as well as the discontinuity and extreme points of the second order partial derivative $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of the photon Hopfield coefficients marked by the vertical fine lines in Fig. 2 and 3. In consequence, $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ can track the change of $p_{\rm c}$ with $\zeta$. Comparing with Fig. 6, one can discover that the presence of the cavity breaks the symmetry of $I_{\rm i,o}$ with respect to $\zeta=0$ for configurations HO ST, HE ST, HE TT, and HO-HE ST [Fig. 7(a3,4)]. Moreover, the exciton transport is suppressed by the cavity in the range of $\zeta\neq 0$. For example, $p_{\rm c}$ is high near $\zeta=-0.3$ [Fig. 7(a5)], suppressing the exciton currents in the main channel with the multichains for HO ST [Fig. 7(a3,4)]. On the other hand, polariton-assisted transport boosts $I_{\rm i,o}$ and $\eta$ near $\zeta=0$ for configurations HE ST, HE TT, and HO-HE TT, in which $p_{\rm c}$ is high, leading the increase of exciton occupation number $p_{N}$ on the last column sites [Fig. 7(a4)]. In fact, $\zeta=0$ means that only one chain undertakes the exciton transport, in which polariton-assisted exciton transport is demonstrated wei2022cavity . Because of few excitations injected [Fig. 7(a1)] for HO ST, HO TT and few $p_{N}$ [Fig. 7(a4)] for HO-HE ST near $\zeta=0$, the cavity has no effect on exciton transport. What’s more, it can be seen obviously that the peaks of $p_{\rm c}$ and of $I_{\rm i,o}$ are at different locations. The reason is that at the peak of $p_{\rm c}$ the transfer of excitation in the multichains to the cavity suppresses the exciton transport that the multichains server as dominating channel, then causing relatively low $I_{\rm i,o}$. Adding the cavity drive, $p_{\rm c}$ increase dramatically but without proportional $I_{\rm i,o}$ and $\eta$ enhancement as shown in Fig. 7(b5). Because of the transfer of excitations, $p_{11}$ increases near $\zeta=0$ [Fig. 7(b1)], causing $p_{\rm c}$ decrease for HE ST, HE TT, HO-HE ST, and HO-HE TT [Fig. 7(b5)], and suppressing the exciton transport of the main channel with the porlaritons [Fig. 7(b3,4,6)]. However, one can see by comparing Fig. 7(b3,4) and Fig. 7(a3,4) that the increase of $p_{\rm c}$ suppresses significantly the exciton currents of the main channel with the multichains near $\zeta=0.2$ for HO-HE ST and near $\zeta=0.35$ for HO ST. Therefore, the exciton transport can be adjusted by the cavity photon. For different configurations, we have the conclusion that ST coupling outperforms TT for the exciton transport. Besides, the optimal $\zeta$ ranges differ for HO and HE chains, which are consistent with the spectral features. In a word, the photon-exciton competition governs the suppression and enhancement of exciton transport via cavity.
Fig. 8(a1-6) and 8(b1-6) show the dependence of the steady-state dynamics on the dimerization parameter $\delta$ under the TLS and the TLS+cavity drive in the presence of cavity, respectively. We analyse firstly the effects of $\delta$ and cavity on exciton transport under the TLS drive. Although the exciton transport efficiencies are centered approximatively on $\delta=0$ [Fig. 8(a6)], there are no the exciton currents in $\delta<0$ except for the configurations HE ST and HO-HE ST of $N=9$ [Fig. 8(a3,4)]. The root cause is the blocked excitation injection at the site (1,1) [Fig. 8(a1,3)]. It can be seen that for the configurations HE ST and HO-HE ST of $N=8$ in the range $0.350<\delta<$ 0.35 the presence of cavity suppresses significantly the exciton currents $I_{\rm i,o}$ [Fig. 8(a3,4)] and transport efficiency $\eta$ [Fig. 8 (a6)] by comparing with Fig. 6(b3-5). Particularly, at $\delta=0.26$ for HE ST and $\delta=0.36$ for HO-HE ST of $N=8$, the troughs of $I_{\rm i,o}$ and $\eta$ appear, which correspond to the peak and the trough of $p_{\rm c}$ [Fig. 8(a5)] and to the dressed nearly-zero-energy crossings [Fig. 4], respectively. The reason for the former is that the excitations on the edge-column TLSs are transferred to cavity through coherent exciting at the crossings as shown in Fig. 8(a1,3,4), suppressing the exciton currents and transport efficiency on the main channel with multichains. But for the latter, the occupation number $p_{N}$ is too few, limiting the exciton transport. At $\delta>0.26$ for HE ST and $\delta>0.36$ for HO-HE ST of $N=8$, the limited photon number $p_{\rm c}$ [Fig. 8(a5)] and the decreased occupation number $p_{\rm bulk}$ [Fig. 8 (a2)] lead to that the polariton becomes the main channel, thus the exciton transport is strengthened by the cavity comparing with Fig. 6(b3,4). However, for even-length HO-HE TT both the multichains and the polariton dominate the exciton transport because of the large $p_{\rm bulk}$ and $p_{\rm c}$. For odd-length HO-HE TT, in $\delta\gtrsim 0.35$ the low $p_{N}$ is main reason limiting exciton transport. When the cavity drive is added based on the TLS drive, there are distinct effects on exciton transport in $0.350<\delta<$ 0.35 for HE ST and HO-HE ST of $N=8$. The increased $p_{\rm c}$ [Fig. 8(b5)] leads to more polaritons, enhancing the polariton channel [Fig. 8(b3,4)] than that under TLS drive [Fig. 8(a3,4)], but in $\delta>0.35$ because of the low $p_{\rm c}$ and $p_{\rm bulk}$ [Fig. 8(b2)] the exciton transport is limited for even-length HE ST. On the contrary, because of the large $p_{\rm c}$ and $p_{\rm bulk}$ the exciton transport is heightened for HO-HE ST and HO-HE TT with odd length, as well as for HO-HE ST in $\delta>0.7$ compared with that under TLS drive. Particularly, the exciton transport currents and efficiency are significantly large due to the cavity drive in $\delta<0$ for all configurations of $N=9$, which break the conventional preference that the transport are superior in $\delta>0$ than that in $\delta<0$. In the condition, the high $p_{\rm bulk}$ [Fig. 8(b2)] and $p_{\rm c}$ [Fig. 8(b5)] demonstrate that both the multichains and the polaritons participate the exciton transport. Interestingly, in the presence of the cavity drive $p_{11}$ is no longer necessary to transport efficiently exciton [Fig. 8(b1)], but $p_{N}$ still is. For instance, the low $p_{N}$ limits the exciton transport for HE TT and HO-HE TT of $N=9$ in $\delta<0$. To sum up, in $\delta>0$ the cavity have significant influence on exciton transport for even-length multichains under TLS drive, which is suppressed in weak $\delta$ and is enhanced in strong $\delta$. In the space of $\delta$, even-length ST is more beneficial to exciton transport than TT under TLS drive, but odd-length ST is under TLS+cavity drive. In addition, no matter which one is main for multichain and polariton channels, the occupation of terminal sites decides the exciton transport.
When the TLS transition frequency detuning $\Delta$ is considered, we plot the steady-state dynamics of the hybrid system in Fig. 9 and 10, which show respectively the effects of the inter-chain coupling $\zeta$ and the dimerization parameter $\delta$ for different configurations. Comparing with Fig. 2 - 5, the effects of $\zeta$, $\delta$, and $\Delta$ on $\eta$ mirror spectral structures.
In Fig. 9 the zones of $\zeta$ and $\Delta$ that the cavity has obviously effects on exciton transport are marked in Fig. 9(ab-fb, g-l, and gc-lc) by ellipses through comparing to Fig. 9(a-f). According to the effects these coupling configurations of the multichains can be classified into two categories. One is HO ST and HO TT that the exciton transport is suppressed, the others has opposite influence. For the former, the excitation on whole multichains transfer to the cavity disrupts the multichains transport serving as dominating channel [Fig. 9(gc,hc)]. For the letter, as shown in Fig. 9(cb-fb) too small $p_{\rm bulk}$ limits the multichain channel in the marked zone, but the large $p_{\rm c}$ strengthens the polariton channel [Fig. 9(ic-lc)]. We can further illustrate the mechanism through the spectra. Interestingly, the structures of $p_{\rm c}$ [Fig. 9(gc-lc)] are similar to those of the cavity-dressed energy levels [Fig. 2 and 3]. In the intensified zone by the cavity the cavity-dressed anticrossings exist [Fig. 2 and 3], thus LZ transition leads to that the excitations on multichains generate polaritons. Particularly, in the zone with high efficiency there are two anti-crossed zonate zones with large $p_{\rm c}$ [Fig. 9(ic)] corresponding to the cavity-dressed anticrossings for HE ST. When the cavity drive is exerted, for the two categories of configurations the exciton transport is further suppressed and enhanced by the cavity respectively. HE ST shows the biggest effect of cavity drive on exciton transport in six configurations, and that the high efficient zone is significantly extended.
In parameter space of $\delta$ and $\Delta$, as shown in Fig. 10 the steady-state dynamics depend on the coupling configuration, the cavity, and the method of drive. For HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ the cavity has apparent effects on exciton transport efficiency in the zone of the fully dimerized limit, namely, $\delta\rightarrow\pm 1$, and near the zero detuning [Fig. 10(a-d, g-i, m-p)]. Nevertheless, it is contrary for HO-HE TT [Fig. 10(s-v)] because the HO chain and the diagonal inter-chain coupling suppress the dimerized effect. Undoubtedly, in the zone enhancing exciton transport the polaritons act as prime role whose mechanism is coherent excitation (LZ transition) located at the cavity-dressed crossings (anticrossings) according to the analysis on Fig. 4, 5 and 8. When the cavity drive is added, it can be seen that the high efficient zones are extended and the exciton transport efficiency is strengthened in the zone of $\delta\rightarrow\pm 1$ and near $\Delta=0$ for HE ST, HE TT, and HO-HE ST of $N=8,~{}9$ [Fig. 10(e, f, k, l, q, r)]. In point of the optimal exciton transport, HE ST is the preferred configuration in the parameter space of $\zeta$ and $\Delta$ under the TLS drive in the presence of the cavity, as well as HE TT is the preferred configuration in the parameter space of $\delta$ and $\Delta$ under the TLS+cavity drive. Significantly, according to the results the exciton transport can be controlled by the coupling configurations, the cavity, the drive methods, the inter-chain coupling, the dimerization parameter, and the TLS transition frequency detuning.
IV.2 The effects of the number and the length of chains
In this part, we discuss the dependence of exciton transport dynamics on the chain number $L$ and length $N$, which reveal the inter-chain joint effect, and give optimal $L$, $N$, and configurations with high transport efficiency and currents in Fig. 11 and 12.
Fig. 11(c1111, g1111, d1111) shows that in the absence of the cavity the exciton transport is mightily dependent on chain number $L$ for the three configurations HE ST, HO-HE ST, and HO-HE TT when $L\leq 6$, and is opposite for the others. For HE TT, although the input exciton current $I_{i}$ is large, the little $p_{N}$ limits the outgoing exciton current $I_{o}$, leading to that exciton transport is independent of $L$. In fact, the inefficient exciton transport among chains lead to excitons localizing at the upper edge sites [Fig. 11(b1)], which is similar to skin effect. Contrary to the skin effect, for HE ST because of the diffusion of excitons among chains the inner occupation number $p_{\rm bulk-in}$ increase and that of the edge decrease with $L$ as shown in Fig. 11(b1, e1, f1). For large $L$ the excitons collect at the inside of the multichains [Fig. 11(e1)], thus HE ST is a potential candidate applied to quantum battery in the six configurations. In the presence of the cavity the steady-state dynamics present characteristic phenomena as shown in Fig. 11(a2-h2). The prominent one is that the exciton currents and transport efficiency present distinctly odd-even oscillation with $L$ for HE TT and HO-HE TT, in which the peaks (troughs) of $I_{\rm i,o}$ and $\eta$ correspond to the odd (even) $L$ when $N=10$ [Fig. 11(c2, g2, d2)]. The difference between the steady-state dynamics of the odd and even $L$ becomes small with $L$. In addition, it can be seen that only for HE ST and HE TT the exciton transport is strengthened by the cavity when $N=10$.
There is an obvious character in Fig. 12 that the exciton transport efficiency $\eta$ also behaves odd-even oscillation with the chain length $N$. When the cavity is absent, the oscillation is weak for HO ST relative to the other configurations and depends simultaneously on the odevity of $L$ and $N$ for HO TT [Fig. 12(a-d)]. Thus the diagonal inter-chain coupling induces the difference between the exciton transports of odd and even $N$ in long HO TT multichains as shown in Fig. 12(b, d). It can be found that because of the dimerization the prominent oscillation with $N$ exists in the short multichains, but disappears in the long multichains for HE ST and HE TT [Fig. 12(a-d)]. In fact, the dimerization hinders the exciton diffusing on the multichains for the later. However, in the mixed long multichains of HO-HE, because HO chains suppress the dimerization, the oscillation is enhanced for HO-HE ST and HO-HE TT. For the two configurations the oscillations with $N$ are consistent for odd $L$ and are opposite for even $L$ each other, such as when $L=3$ and $N=16$ $\eta$ is trough for both HO-HE TT and HO-HE ST, while $L=4$ and $N=16$ $\eta$ is peak for HO-HE ST and is trough for HO-HE TT. When the cavity is considered the oscillation is significantly strengthened for the configurations containing HE chain, as shown in Fig. 12(e-h). For HE ST the oscillation decreases with $L$, which is caused by the dissipation of the more and more TLSs. Whatever situation the oscillation of $\eta$ with $N$ for HE ST and HE TT is consistent, which illustrates further that the oscillation is mainly caused by HE chain. Interestingly, the oscillation amplitude behaves inverse dependence on the odevity of $L$ between HE TT and HO-HE ST, HO-HE TT. Concretely, for the former the amplitude with odd $L$ is bigger than that with even $L$, and for the latter it is contrary. Furthermore, it can be discovered that for HO-HE ST and HO-HE TT the odd-even oscillation of $\eta$ with $N$ can be adjusted by $L$ and the cavity. For example, the oscillations of $L=3$ and $L=4$ for HO-HE ST are out of phase [Fig. 12(a, b)]. $\eta$ without [Fig. 12(b)] and with [Fig. 12(f)] cavity for HO-HE ST of $N=14$ and $L=2$ are respectively trough and peak. Comparing Fig. 12(b) and 12(f), Fig. 12(f) and 12(g) for HO-HE ST, the period of the oscillation also is varied by the cavity and $L$.
V Summary
In summary, in order to reveal the exciton transport mechanism of a 2D multichains system within a cavity, we investigate systematically the spectra through numerically diagonalizing the Hamiltonian, the second order partial derivative of photon Hopfield coefficients with respect to the inter-chain coupling strength and the dimerization parameter, the von Neumann entropy and the steady-state dynamics of exciton via solving the Lindblad quantum master equation, in which six inter-TLS coupling configurations are considered as HO ST, HO TT, HE ST, HE TT, HO-HE ST and HO-HE TT based on the HO and HE chain, as well as the square and triangular inter-chain coupling type. The results show that the exciton transport can be adjusted by the inter-chain coupling strength, the dimerization parameter, the detuning of the driving frequency, the chain number and length, the coupling configurations, and the cavity. In the absence of cavity the exciton transport is decided by the distribution of exciton on whole multichains, in which the efficient transport need the excitons spread all over the multichains. Further, when the cavity is considered it is governed by the competition between the polariton and multichain channel. When the excitations are mainly distributed on whole multichains, the multichains dominate the transport. Otherwise, the cavity occupies the most excitations, the polaritons act as the main channel, meanwhile the limited exciton occupations of the multichain terminals are necessary. The exchange condition of the excitation between the multichains and cavity is the key which channel is primary. The number of polaritons depend on the photon occupation number, which is relate to the cavity-dressed nearly-zero-energy crossings and anticrossings. At the crossings and anticrossings the coherent excitation and LZ transition between the multichains and cavity occur, respectively. Therefore, the exciton transport is controlled by the crossings and anticrossings, whose locations can be accurately determined by von Neumann entropy. Besides, we discover that the second order partial derivative of the photon Hopfield coefficient serves as robust indicator for identifying the crossings and anticrossings, and the change of photon occupation number. The spectra show that the crossings and anticrossings are related closely to the coupling configurations, the inter-chain coupling strength, and the dimerization parameter. In the space of inter-chain coupling strength, ST outperforms TT for the exciton transport, but the optimal ranges differ for HO and HE chains. In the space of the dimerization parameter, even-length ST is more beneficial to exciton transport than TT under TLS drive, but odd-length ST is under TLS+cavity drive. Interestingly, when the cavity drive is exerted, the exciton transport is not limited by the occupation number of site (1,1) and is strengthened for HE ST and HO-HE ST with odd chains in the range of the negative dimerization parameter. Further, we discover that the steady-state dynamics can be adjusted by the TLS transition frequency detuning. Finally, it can be found that the exciton transport currents and efficiency present distinctly odd-even oscillation with chain length and number, which depends on the coupling configuration and the cavity. In a word, our results reveal the exciton transport mechanism of the 2D hybrid system, which provides theoretical foundation for designing controllable and high efficient exciton transport devices, such as solar cell. Of course, it has many points of view to reveal the mechanism, like the topological phase transition of the system.
VI Acknowledgements
This work is supported by the Academic Ability Promotion Foundation for Young Scholars of Northwest Normal University in China under the Grant No. NWNU-LKQN2024-10, National Natural Science Foundation of China under Grant No. 12164042, National Natural Science Foundation of China under Grant No. 12104374, National Natural Science Foundation of China under Grant No. 12065022.